### v4

# Multi-parameter Module Approximation: an efficient and interpretable invariant for multi-parameter persistence modules with guarantees

Abstract
Topological data analysis (TDA) is a rapidly growing area of data science, whose most common descriptor is persistent homology, which tracks the topological changes in growing families of subsets of the data set itself, called filtrations, and encodes them in an algebraic object, called a persistence module. The algorithmic and theoretical properties of persistence modules are now well understood in the single-parameter case, that is, when there is only one filtration (e.g., feature scale) to study. In contrast, much less is known in the multi-parameter case, where several filtrations (e.g., scale and density) are used simultaneously. Since multi-parameter persistence modules usually encode information that is invisible to their single-parameter counterparts, it is critical to build tractable proxies for them, ideally with some theoretical robustness guarantees.
In this article, we introduce a new parameterized family of topological descriptors, taking the form of candidate decompositions, for multi-parameter persistence modules, and we a identify a subfamily of these descriptors, that we call approximate decompositions, that are controllable approximations, in the sense that they preserve diagonal barcodes. Then, we introduce MMA (Multipersistence Module Approximation): an algorithm based on matching functions for computing instances of candidate decompositions with some precision parameter $\delta>0$. By design, MMA can handle an arbitrary number of filtrations, and has bounded complexity and running time. Moreover, we prove the robustess of MMA: when computed with so-called compatible matching functions, we show that MMA produces approximate decompositions (and we prove that such matching functions exist for $n=2$ filtrations). Next, we restrict the focus on modules that can be decomposed into interval summands. In that case, compatible matching functions always exist, and we show that, for small enough $\delta$, the approximate decompositions obtained with such compatible matching functions by MMA have an approximation error (in terms of the standard interleaving and bottleneck distances) that is bounded by $\delta$, and that reaches zero for an even smaller, positive precision $\delta_{\rm exact}$. Finally, we present empirical evidence validating that MMA has state-of-the-art performance and running time on several data sets.
Keywords: Multiparameter Persistent Homology, Persistence Modules, Interval Modules, Approximation Methods, Convergence Analysis
1 Introduction
Topological Data Analysis (TDA) (35, 54) is a new and rapidly developing area of data science that has seen a lot of interest due to its success in various applications, ranging from bioinformatics (55) to material science (16). The main computational tool of TDA is persistent homology (PH). Whereas homology is a qualitative descriptor of the shape of a topological space $S$, the core idea of PH is to capture how the homology groups change when computed on a filtration of $S$. A filtration is a family $\{S_{x}\subseteq S\}_{x\in A}$ of subspaces of $S$ indexed over a partially ordered set (poset) $A$, that is nested w.r.t. inclusion, i.e., it satisfies $S_{x}\subseteq S_{y}$ for any $x\leq y$. Then, the functoriality of homology and these inclusions induce morphisms between the corresponding homology groups $H_{*}(S_{x})\rightarrow H_{*}(S_{y})$ for each pair $x\leq y$, which allows to detect the differences in homology when going from index $x$ to index $y$. One of the most common ways to produce such filtrations is to study the sublevel sets of a continuous filter function $f:S\rightarrow\mathbb{R}^{n}$, defined with $S_{x}=f^{-1}(\{x^{\prime}\in\mathbb{R}^{n}:x^{\prime}\leq_{n}x\})$; the partial order on the poset $\mathbb{R}^{n}$ (denoted by $\leq_{n}$) is defined, for $x,y\in\mathbb{R}^{n}$, as $x\leq_{n}y$ if and only if $x_{i}\leq y_{i}$ for every dimension $i$.
Single-parameter PH.
When $\mathcal{A}$ is totally ordered, e.g., when $\mathcal{A}\subseteq\mathbb{R}$, then applying the homology functor $H_{*}(-;k)$ for a field $k$ to a (single-parameter) filtration results in a sequence of vector spaces connected by linear maps, called a single-parameter persistence module. This situation has been studied extensively in the TDA literature (18, 24, 35, 54). Notably, one can show that such persistence modules can always be decomposed into a direct sum of simple interval summands: ${\mathbb{M}}\simeq\bigoplus_{i\in\mathcal{I}}{k^{I(b_{i},d_{i})}}$, where each interval summand $k^{I(b_{i},d_{i})}$ intuitively represents the lifetime $I(b_{i},d_{i})$ of a topological feature, i.e., $b_{i}$ is the appearance time (birth) and $d_{i}$ is the disappearance time (death) of a topological feature, that is detected by homology as the index increases. Moreover, single-parameter persistence modules can be efficiently represented in a compact descriptor called the persistence barcode, and several representation methods, such as Euclidean embeddings and kernels for machine learning classifiers, have been proposed for such barcodes in the literature (15, 1, 56, 22, 21). As a consequence, most applications of TDA use single-parameter persistence modules, and often use the sublevel sets of, e.g., the data set scale, as the corresponding single-parameter filtration.
Multi-parameter PH.
However, many data sets come with not just one, but multiple, possibly intertwined, salient filtrations. For example, image data typically has both a spatial filtration and an intensity filtration, and arbitrary point cloud data can be filtered both by feature scale and density. Unfortunately, in general, the resulting multi-parameter persistence modules, obtained by applying the homology functor to a filtration indexed over $\mathbb{R}^{n}$ (13), are much less tractable; in contrast to the single-parameter case, there is no decomposition theorem that can break down any module into a direct sum of simple indicator summands (e.g., interval modules). Instead, there is now a rich literature on general decompositions into arbitrarily complicated summands (34, 32) and their associated minimal presentations (36, 38, 48), on the theoretical study of a few restricted cases (such as some specific $n=2$-parameter filtrations or exact p.f.d. (Definition 2) $2$-parameter persistence modules) where simple decompositions can be obtained (27, 10, 4, 11, 44), and on simpler representations of multi-parameter persistence modules, such as the Euler characteristic (37), the fibered barcode (29, 57), the signed barcode (14, 51), and the (generalized) rank invariant and persistence diagram (41, 59). It has thus become crucial to define general topological descriptors for multi-parameter persistence modules that are meaningful, visually interpretable, and easily computable.
Contributions.
In this article, we introduce new descriptors for multi-parameter persistence modules ( computable from multi-parameter filtrations of simplicial complexes [^2] Or even solely from module presentations, see Remark 13.) along with a new algorithm that we call MMA (Multipersistence Module Approximation) for their practical computations.
[…]
  1. 1.
    We fix a grid of diagonal lines $L$ spaced by $\delta>0$, and compute the barcodes associated to ${({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}})}_{l\in L}$.
  2. 2.
    Using some matching function $\sigma$, we match bars of consecutive barcodes together.
  3. 3.
[…]
Our contributions are five-fold:
  1. 1.
    We introduce a new family of topological descriptors for finitely presented multi-parameter persistence modules (Definition 19), taking the form of candidate decompositions $\tilde{\mathbb{M}}_{\delta}=\bigoplus_{i\in\tilde{\mathcal{I}}}{k^{\tilde{I}_{i}}}$. These candidate decompositions are parameterized by a precision parameter $\delta>0$, and each $k^{\tilde{I}_{i}}$} in these candidate decompositions is an interval summand in $\mathbb{R}^{n}$.
  2. 2.
    Then, we introduce our method MMA (Multi-parameter persistence Module Approximation, Algorithm 1) for computing instances of such candidate decompositions. Our method is crucially based on so-called matching functions, and, using any matching function whose complexity is linear w.r.t. $N$, has running time
    $O\left(N^{3}+\frac{1}{\delta^{n-1}}(N+n\cdot 2^{n-1})\right),$
    where $N$ is the number of simplices and $n$ is the number of filtrations. See Figure 2. Note that MMA does not require the input module $\mathbb{M}$ to be interval decomposable in order to run.
  3. 3.
    We show that MMA is a good approximation when computed with so-called compatible matching functions, i.e., that the candidate decompositions produced by MMA are, in this case, approximate decompositions: they preserve the (single-parameter) persistence barcodes associated to diagonal slices of the multi-parameter filtration (Proposition 3):
    $\textnormal{for all diagonal line }l\subseteq\mathbb{R}^{n},\quad{d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}\vphantom{\big|}\right|_{l}}\right)\right)\leq 2\delta.$
    We also show that, upon carefully choosing matching functions, the approximate decompositions produced by MMA are also stable w.r.t. the input data (Proposition 4):
    ${d_{\rm I}}(\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}(f),\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}(g))\leq{d_{\rm B}}(\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}(f),\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}(g))\leq\left\lVert f-g\right\rVert_{\infty}+\delta,$
    where $\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}(f)$ stands for the candidate decomposition that is induced by the sublevel sets of $f$ and computed with MMA (and similarly for $g$).
  4. 4.
    Then, we restrict the focus to interval decomposable modules. In that case, compatible matching functions always exist, and, under generic assumptions and small enough $\delta$, we prove that the interleaving and bottleneck distances between the approximate decompositions produced by MMA and the underlying persistence module are upper bounded (Proposition 5):
    ${d_{\rm I}}(\mathbb{M},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})\leq{d_{\rm B}}(\mathbb{M},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})\leq\delta.$
    Moreover, we also show that, when $\delta\leq\delta_{\rm exact}$, where $\delta_{\rm exact}$ is a constant that depends only on the multi-parameter filtration values, the approximate decompositions produced by MMA recover the underlying persistence module exactly (Proposition 12):
    ${d_{\rm I}}(\mathbb{M},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})={d_{\rm B}}(\mathbb{M},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})=0.$
  5. 5.
    We perform numerical experiments that showcase the performance of MMA and exhibit the trade-off between computation time and approximation error in Section 7.
Figure 2: Example of candidate decomposition computed by MMA on a point cloud filtered by both growing balls around the points (also called Čech filtration) and using the sublevel sets of codensity (or, equivalently, the superlevel sets of density), in homology dimension 1. One can see that there is a large lightgreen summand in the candidate decomposition on the right that corresponds to the cycle formed by the points amid outliers, which is also highlighted in lightgreen in the multi-parameter filtration on the left.
Related work.
Computing approximate decompositions of multi-parameter persistence modules with simple summands or interval summands has been studied in a few other works. For instance, in (33), the authors provide an algorithm that computes optimal approximations with rectangle summands of interval decomposable, $2$-parameter persistence modules in order to lower bound their interleaving distance. In (5), the authors provide a method that associates to every $2$-parameter persistence module (and that was recently generalized to any $n$-parameter persistence module (6)) a pair of interval decomposable persistence modules obtained by computing the Möbius inversion of the so-called compressed multiplicity, in a spirit similar to the computation of generalized persistence diagrams and signed barcodes. Focusing on homology dimension 0 and $n=2$ with the Vietoris-Rips bifiltration on augmented metric spaces, the authors in (17) have proposed the elder-rule staircode, an interval decomposable persistence module who is known to recover the rank invariant and even the true interval decomposition of the module (if it exists) under some assumptions. Focusing rather on stability (as general decompositions of modules are known to be highly instable), in (7) the author provides new decompositions for persistence modules based on pruning (although with arbitrarily complicated summands) that enjoy better robustness guarantees.
More closely related to our approach, both (20) and (39, 40) provide methods to compute descriptors for $2$-parameter persistence modules using matching functions between persistence barcodes computed from a sorted family of $1$-dimensional slices, namely the multi-parameter persistence image, and the graphcode, respectively. While the multi-parameter persistence image encodes a decomposition constructed by matching consecutive barcodes with the vineyards algorithm (28) and computing the summand boundaries with the endpoints of matched bars (which does not guarantee that the resulting summand is an interval, or even that the resulting module is close to the input), the graphcode is an abstract graph whose nodes represent the bars of consecutive barcodes, and whose edges are built based on matching functions representing canonical inclusions of representative cycles, after fixing some cycle bases for every slice. Note that these matchings are not one-to-one: a representative cycle in one barcode might be included in several others in the next barcode.
Our method differs from the previous ones in three key aspects: first, note that, except for (6, 7), all other approaches are designed for $n=2$ filtrations, while our method MMA can handle an arbitrary number $n$ of filtrations (albeit without theoretical guarantees if $n>2$) and works in any homology dimension. Second, we designed MMA so that it has parameterized complexity: while the other approaches might be costly to compute as they rely on, e.g., Möbius inversions on large non-grid posets, the running time for MMA is controlled by the user through the choice of $\delta$ and of the matching function. Finally, and most importantly, our intention with the interval decomposable module produced by MMA was to provide a descriptor that is:
  1. (i)
    interpretable in the same way than persistence barcodes: its summands should correspond to lifetimes (in $\mathbb{R}^{n}$) of some homologous representative cycles, and
  2. (ii)
    as stable as possible: as it is impossible to provide interval decompositions that are always consistent with the rank invariant of the module (see (46, Section 10.2.3)), and as we still want to encode more information than the pointwise dimension of the module (i.e., the Hilbert function, which already requires a cubic complexity to compute in the two parameter case (26)), we seek for decompositions that preserve the persistence barcodes of diagonal slices of the module.
We prove in this article that our method MMA produces interval decompositions that satisfy both Items (i) and (ii). On the other hand, other approaches are either less interpretable (33, 5, 6, 7, 39, 40), thus not satisfying (i), or not stable enough (20, 17), thus not satisfying (ii).
More precisely, concerning (ii), our method can be seen as a generalization of the decompositions provided in (17), as we are not restricted to homology dimension 0, and as a continuation of the decompositions provided in (20), as $(a)$ we generalize it to compatible matching functions instead of relying solely on the vineyards algorithm, $(b)$ we guarantee that the produced decompositions only contain interval summands instead of producing summands supported on arbitrarily complicated shapes, and $(c)$ we prove that the produced decomposition preserve the diagonal barcodes, and recover the true decompositions exactly when the input module is itself interval decomposable.
Finally, one might argue that the interpretability property that we ask in (i) also induces some degree of arbitrariness, as there are, in general, several ways to choose representative cycles and their lifetimes (precisely because there are no good barcodes in multi-paramerer persistent homology, again see (46, Section 10.2.3)). This is perfectly true, and illustrates the trade-off between interpretability and computability that is often encountered in this field; note for instance that graphcodes (39, 40) also rely on arbitrary choices of cycles bases. However, notice that $(a)$ using such interval decompositions has nonetheless already proved to be useful in computational topology (8) and data science (49), and $(b)$ the collection of interval decompositions produced by MMA obtained by ranging over all compatible matching functions is itself a topological invariant of the module.
Outline.
Section 2 provides a concise review of multi-parameter persistence modules and related notions. In Section 3, we present our new descriptors for multi-parameter persistence modules, as well as the MMA algorithm for computing them. In Section 4, we present some approximation properties satisfied by MMA. Then, we provide stronger robustness guarantees for interval decomposable modules in Section 5, and we discuss the design of matching functions in Section 6. Finally, we illustrate the performances of MMA in Section 7.
2 Background
In this section, we recall the basics of multi-parameter persistent homology and persistence modules. This section only contains the necessary background and notations, and can be skipped if the reader is already familiar with persistence theory. A more complete treatment of persistence modules can be found in (24, 31, 54, 13).
Notations.
We first introduce a few notations: we let $(e_{1},\dots,e_{n})$ be the canonical basis of $\mathbb{R}^{n}$, and, given a set $A\subseteq\mathbb{R}^{n}$, we let ${\rm conv}(A)$ denote the convex hull of $A$. Moreover, given a hyperplane $H\subseteq\mathbb{R}^{n}$ and its two associated vectors $a_{H},b_{H}\in\mathbb{R}^{n}$ which satisfy $H=b_{H}+\{x\in\mathbb{R}^{n}:\langle x,a_{H}\rangle=0\}$, we call $a_{H}$ the codirection of $H$. When $a_{H}$ is a vector in the canonical basis of $\mathbb{R}^{n}$, i.e., there exists $i\in\llbracket 1,n\rrbracket$ such that $a_{H}=e_{i}$, we slightly abuse notation and also call $i$ the codirection of $H$. Finally, given a point $x\in\mathbb{R}^{n}$, we let $\left<x\right$< (resp. $\left>x\right$>) denote the upset (resp. downset) of $x$: $\left<x\right<:=\left\{y\in\mathbb{R}^{n}:y\geq_{n}x\right\}$ (resp. $\left>x\right>:=\left\{y\in\mathbb{R}^{n}:y\leq_{n}x\right\}$).
Multi-parameter persistence modules.
In their most general form, multi-parameter persistence modules (19) are nothing but $k$-vector spaces (where $k$ denotes a field) indexed by $\mathbb{R}^{n}$ and connected by linear maps.
Definition 1 (Multi-parameter persistence module).
An ($n$-)multi-parameter persistence module $\mathbb{M}$ is a family of vector spaces indexed over $\mathbb{R}^{n}$: $\mathbb{M}=\{M_{x}\}_{x\in\mathbb{R}^{n}}$, equipped with linear transformations $\{\varphi_{x}^{y}:M_{x}\rightarrow M_{y}\}_{x,y\in\mathbb{R}^{n},x\leq_{n}y}$, that are called the transition maps of $\mathbb{M}$, and that satisfy $\varphi_{x}^{z}=\varphi_{y}^{z}\circ\varphi_{x}^{y}$ and $\varphi^{x}_{x}=\mathrm{id}$ for any $x\leq_{n}y\leq_{n}z$. We sometimes let $\mathbb{M}(x\leq_{n}y):=\varphi_{x}^{y}$ denote these transition maps.
A morphism between two multi-parameter persistence modules $\mathbb{M},\mathbb{M}^{\prime}$ with transition maps $\varphi^{\cdot}_{\cdot}$ and $\psi^{\cdot}_{\cdot}$ respectively, is a collection of linear maps $f=\{f_{x}:M_{x}\rightarrow M^{\prime}_{x}\}_{x\in\mathbb{R}^{n}}$, that commutes with transitions maps, i.e., one has $f_{y}\circ\varphi^{y}_{x}=\psi^{y}_{x}\circ f_{x}$, for all $x\leq_{n}y$.
Multi-parameter persistence modules are often assumed to satisfy finiteness assumptions, such as being pointwise finite dimensional (Definition 2) or finitely presentable (Definition 9). See (13) for more details.
Definition 2 (Pointwise finite dimensional module).
Let $\mathbb{M}$ be an $n$-parameter persistence module. We say that $\mathbb{M}$ is pointwise finite dimensional (p.f.d.), if for any $x\in\mathbb{R}^{n}$, one has $\dim\mathbb{M}_{x}<\infty$.
In this article, all multi-parameter persistence modules come from applying the homology functor $H_{*}$ on a multi-parameter filtration of a topological space $S$, that is, on a family $\{S_{x}\}_{x\in\mathbb{R}^{n}}$ of subsets of $S$ indexed over $\mathbb{R}^{n}$ such that $x\leq_{n}y\Rightarrow S_{x}\subseteq S_{y}$. In other words, we study modules of the form $\mathbb{M}:=\{H_{*}(S_{x})\}_{x\in\mathbb{R}^{n}}$, where the linear maps $H_{*}(S_{x})\to H_{*}(S_{y})$ are induced by the canonical inclusions $S_{x}\subseteq S_{y}$ (when $x\leq_{n}y$). There are many interesting multi-parameter filtrations in data science; one of the most common one (with $n=2$) comes from filtering by feature scale and density. This allows to detect the topological structures (encoded in the homology groups) of point clouds in the face of noise and outliers (9, 20).
Definition 3.
The direct sum of two multi-parameter persistence modules $\mathbb{M}$ and $\mathbb{M}^{\prime}$, written as $\mathbb{M}\oplus\mathbb{M}^{\prime}$, is the module $\mathbb{M}^{\prime\prime}$ with vector spaces $\{M^{\prime\prime}_{x}\}_{x\in\mathbb{R}^{n}}$ and transition maps $(\varphi^{\prime\prime})_{\cdot}^{\cdot}$, defined as $M^{\prime\prime}_{x}=M_{x}\oplus M^{\prime}_{x}$ for all $x\in\mathbb{R}^{n}$, and $(\varphi^{\prime\prime})_{\cdot}^{\cdot}=\varphi_{\cdot}^{\cdot}\oplus(\varphi^{\prime})_{\cdot}^{\cdot}$, where $\{M_{x}\}_{x\in\mathbb{R}^{n}}$ (resp. $\{M^{\prime}_{x}\}_{x\in\mathbb{R}^{n}}$) and $\varphi_{\cdot}^{\cdot}$ (resp. $(\varphi^{\prime})_{\cdot}^{\cdot}$) are the vector spaces and transition maps of $\mathbb{M}$ (resp. $\mathbb{M}^{\prime}$) respectively.
A multi-parameter persistence module $\mathbb{M}$ such that there are no non-trivial modules $A$ and $B$ such that $\mathbb{M}\simeq A\oplus B$ is called indecomposable.
Note that while multi-parameter persistence modules can always be decomposed into indecomposable summands if they are p.f.d. (see (34, 32) for corresponding algorithms), these summands can be arbitrarily complicated, and the resulting decomposition cannot really be used as an intuitive and simple invariant of the module.
Distances between modules.
Multi-parameter persistence modules can be compared with the standard interleaving distance (45).
Definition 4 (Interleaving distance).
Given $\varepsilon>0$, two multi-parameter persistence modules $\mathbb{M}$ and $\mathbb{M}^{\prime}$ are $\boldsymbol{\varepsilon}$-interleaved if there exist two morphisms $f\colon\mathbb{M}\to\mathbb{M}^{\prime}(\boldsymbol{\varepsilon})$ and $g\colon\mathbb{M}^{\prime}\to\mathbb{M}(\boldsymbol{\varepsilon})$ such that $g_{\cdot+\boldsymbol{\varepsilon}}\circ f_{\cdot}=\varphi_{\cdot}^{\cdot+2\boldsymbol{\varepsilon}}\text{ and }f_{\cdot+\boldsymbol{\varepsilon}}\circ g_{\cdot}=\psi_{\cdot}^{\cdot+2\boldsymbol{\varepsilon}},$ where $\mathbb{M}(\boldsymbol{\varepsilon})$ is the shifted module $\{M_{x+\boldsymbol{\varepsilon}}\}_{x\in\mathbb{R}^{n}}$, $\boldsymbol{\varepsilon}=(\varepsilon,\dots,\varepsilon)\in\mathbb{R}^{n}$, and $\varphi$ and $\psi$ are the transition maps of $\mathbb{M}$ and $\mathbb{M}^{\prime}$ respectively.
The interleaving distance between two multi-parameter persistence modules $\mathbb{M}$ and $\mathbb{M}^{\prime}$ is then defined as ${d_{\rm I}}(\mathbb{M},\mathbb{M}^{\prime}):=\inf\left\{\varepsilon\geq 0:\mathbb{M}\text{ and }\mathbb{M}^{\prime}\text{ are }\boldsymbol{\varepsilon}\text{-interleaved}\right\}$.
The main property of this distance is that it is stable for multi-parameter filtrations that are obtained from the sublevel sets of functions. More precisely, given two continuous functions $f,g:S\to\mathbb{R}^{n}$ defined on a topological space $S$ let $\mathbb{M}(f),\mathbb{M}(g)$ denote the multi-parameter persistence modules obtained from the corresponding multi-parameter filtrations $\{S^{f}_{x}:=\{s\in S:f(s)\leq_{n}x\}\}_{x\in\mathbb{R}^{n}}$ and $\{S_{x}^{g}:=\{s\in S:g(s)\leq_{n}x\}\}_{x\in\mathbb{R}^{n}}$. Then, one has (45, Theorem 5.3):
${d_{\rm I}}(\mathbb{M}(f),\mathbb{M}(g))\leq\left\lVert f-g\right\rVert_{\infty}.$ (1)
Another usual distance is the bottleneck distance (12, Section 2.3). Intuitively, it relies on decompositions of the modules into direct sums of indecomposable summands, and is defined as the largest interleaving distance between summands that are matched under some matching.
Definition 5 (Bottleneck distance).
Given two multisets $A$ and $B$, $\mu\colon A\not\to B$ is called a partial bijection if there exist $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ such that $\mu\colon A^{\prime}\to B^{\prime}$ is a bijection. The subset $A^{\prime}:=\mathrm{coim}(\mu)$ (resp. $B^{\prime}:=\mathrm{im}(\mu)$) is called the coimage (resp. image) of $\mu$.
[…]
The bottleneck distance between two multi-parameter persistence modules $\mathbb{M}$ and $\mathbb{M}^{\prime}$ is then defined as ${d_{\rm B}}(\mathbb{M},\mathbb{M}^{\prime}):=\inf\left\{\varepsilon\geq 0\,:\,\mathbb{M}\text{ and }\mathbb{M}^{\prime}\text{ are }\boldsymbol{\varepsilon}\text{-matched}\right\}.\\$
Since a matching between the decompositions of two multi-parameter persistence modules induces an interleaving between the modules themselves, it follows that ${d_{\rm I}}\leq{d_{\rm B}}$. Note also that $d_{\rm B}$} can actually be arbitrarily larger than $d_{\rm I}$}, as showcased in (12, Section 9).
Interval modules.
Now, we define a particular subfamily of multi-parameter persistence modules, the so-called interval modules. Intuitively, they are modules that are trivial, except on a subset of $\mathbb{R}^{n}$ called an interval.
Definition 6 (Interval).
A subset $I$ of $\mathbb{R}^{n}$ is called an interval if it satisfies:
  • (convexity) if $p,q\in I$ and $p\leq_{n}r\leq_{n}q$ then $r\in I$, and
  • (connectivity) if $p,q\in I$, then there exists a finite sequence $r_{1},r_{2},\dots,r_{m}\in I,$ for some $m\in\mathbb{N}$, such that $p\sim r_{1}\sim r_{2}\sim\dots\sim r_{m}\sim q$, where $\sim$ can be either $\leq_{n}$ or $\geq_{n}$.
Definition 7 (Interval module, indicator module).
An interval module $\mathbb{I}$ is a multi-parameter persistence module such that:
  1. 1.
    $\mathbb{I}$ is a thin module, i.e., $\forall x\in\mathbb{R}^{n},\,{\rm dim}(\mathbb{I}_{x})\leq 1$,
  2. 2.
    whose support ${\rm supp}(\mathbb{I}):=\{x\in\mathbb{R}^{n}:{\rm dim}(\mathbb{I}_{x})=1\}$ is an interval of $\mathbb{R}^{n}$,
  3. 3.
    and whose transition maps are identity maps, i.e., $\forall x\leq_{n}y\in{\rm supp}(\mathbb{I}),\,\mathbb{I}(x\leq y)=\mathrm{id}$.
Moreover, given an interval $I$, we let $k^{I}$ denote the corresponding interval module with support $I$. Finally, a module $\mathbb{I}$ that is a direct sum of interval modules $\mathbb{I}=k^{I_{1}}\oplus\dots\oplus k^{I_{m}}$, $m\in\mathbb{N}^{*}$, whose supports have empty pairwise (closed) intersections, i.e., such that $I_{p}\cap I_{q}=\varnothing$ for all $1\leq p,q\leq m$, $p\neq q$, is called an indicator module, and denoted by $k^{U}$, where $U:=\cup_{i=1}^{m}I_{i}$ is the union of their supports.
Finally, we define a specific type of interval modules, those whose support is equal to a union of rectangles. We call these modules discretely presented—the candidate decompositions computed by our algorithm MMA (Section 3) are actually made up of such modules.
Definition 8 (Discretely presented interval module).
An interval module $\mathbb{I}=k^{I}$ is discretely presented if its support $I$ is a locally finite union of rectangles in $\overline{\mathbb{R}}^{n}$, and whose boundary is an (n1)(n-1)-submanifold of $\mathbb{R}^{n}$. More precisely, there exist two locally finite families of points, the birth and death critical points of $I$, denoted by $C_{B}(I)$ and $C_{D}(I)$ respectively, such that:
$I:=\bigcup_{c\in C_{B}(I)}\bigcup_{c^{\prime}\in C_{D}(I)}R_{c,c^{\prime}},$ (2)
where $R_{c,c^{\prime}}:=\{x\in(\mathbb{R}\cup\{\pm\infty\})^{n}:c\leq_{n}x{{<}_{n}}c^{\prime}\}$ is the rectangle with corners $c$ and $c^{\prime}$.
Discretely presented interval modules can be obtained through stronger assumptions, such as being a finitely presentable interval module. See (13) for more details.
Definition 9 (Finitely presentable persistence module.).
Let $\mathbb{M}$ be an $n$-parameter persistence module. One says that $\mathbb{M}$ is finitely presentable (f.p.) if it is isomorphic to the cokernel of a persistence morphism $f\colon R\to G$, where $G,R$ are interval decomposable modules of the form:
$G=\bigoplus_{1\leq i\leq n_{G}}k^{\left<x_{i}\right<}\quad\textnormal{ and }\quad R=\bigoplus_{1\leq i\leq n_{R}}k^{\left<y_{i}\right<},$ (3)
for some finite sets $\beta_{0}:=(x_{i})_{1\leq i\leq n_{G}}$ and $\beta_{1}:=(y_{i})_{1\leq i\leq n_{R}}$, with $n_{G},n_{r}\in\mathbb{N}$. When $n_{G}$ and $n_{R}$ are minimal, the sets $\beta_{0}(\mathbb{M})$ and $\beta_{1}(\mathbb{M})$ are unique and called the degree 0 and degree 1 graded Betti numbers of $\mathbb{M}$, respectively.
The graded Betti numbers characterize the position of topological events in a multi-parameter filtration. In particular, f.p. modules can be restricted to large enough compact sets without loosing information, if such sets contain the graded Betti numbers, see Remark 1. Next, we define restrictions of modules.
Definition 10 (Restrictions and slices).
If $Q\subseteq\mathbb{R}^{n}$, then an $n$-parameter persistence module $\mathbb{M}=\{M_{x}\}_{x\in\mathbb{R}^{n}}$ induces a persistence module $\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{Q}$} indexed over $Q$, defined by:
$\forall x\leq_{n}y\in Q,\quad\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{Q}}\right)_{x}:=M_{x}\textnormal{ and }{\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{Q}}(x\leq_{n}y):=\mathbb{M}(x\leq_{n}y).$
In particular, if $\mathbb{M}$ is an $n$-parameter persistence module, then given a line $l\subseteq\mathbb{R}^{n}$ with positive slope, the persistence module $\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}$} can be seen as a usual 1-parameter persistence module up to some parametrization of $l$.
Example 1 (Restriction to a line).
[…]
$\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)_{t}:=M_{at+b}\quad\textnormal{and }\quad\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)(s\leq t):=\mathbb{M}(as+b\leq at+b),$ (4)
is a $1$-parameter module, called the restriction, or slice, of $\mathbb{M}$ along $l$.
The opposite operation is called a (left) Kan extension, that we only define in our setup.
Definition 11 (Kan Extension).
Let $A=A_{1}\times\cdots\times A_{n}\subseteq\mathbb{R}^{n}$ be a poset of $\mathbb{R}^{n}$ obtained as a product of subsets of $\mathbb{R}$, and $\mathbb{M}=\{M_{x}\}_{x\in A}$ be a multi-parameter persistence module indexed over $A$. The (left) Kan extensions of $\mathbb{M}$ is the $n$-parameter persistence module (indexed over $\mathbb{R}^{n}$) defined as follows: for any $x\leq_{n}y\in\mathbb{R}^{n}$,
$\left(\mathrm{Lan}_{n}\mathbb{M}\right)_{x}=M_{\lfloor x\rfloor_{A}}\quad\textnormal{and}\quad\mathrm{Lan}_{n}\mathbb{M}(x\leq_{n}y)=\mathbb{M}(\lfloor x\rfloor_{A}\leq_{n}\lfloor y\rfloor_{A}),$ (5)
[…]
Remark 1.
If the graded Betti numbers of $\mathbb{M}$ are included in a rectangle subset $K=[a_{1},b_{1}]\times\cdots\times[a_{n},b_{n}]$ of $\mathbb{R}^{n}$, then the restriction $\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{K}$} contains as much information as the original module $\mathbb{M}$; more formally, we have $\mathrm{Lan}_{n}{\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{K}}\simeq\mathbb{M}$. Notice that, in practice, finding such a compact set $K$ for a given persistence module $\mathbb{M}$ does not require computing a minimal presentation of $\mathbb{M}$ (Definition 9). Indeed, if $\mathbb{M}$ is obtained from applying the homology functor on a finite multi-parameter filtration $F$, i.e., $\mathbb{M}\simeq H_{*}(F)$, then any rectangle $R\subseteq\mathbb{R}^{n}$ containing the filtration values of the simplices of $F$ will also contain the graded Betti numbers.
Note that when only one filtration is given, single-parameter p.f.d. persistence modules (Definition 2) always decompose into interval modules: $\mathbb{M}\cong\bigoplus_{i\in\mathcal{I}}k^{[b_{i},d_{i})}$ (24, Theorem 2.8). In that case, they are frequently represented as the collection of the supports (in $\mathbb{R}$) of their summands, also called persistence barcode ${\mathcal{B}}\left(\mathbb{M}\right):=\{[b_{i},d_{i})\}_{i\in\mathcal{I}}$.
Boundaries and facets of intervals
Now, we recall the definition of upper- and lower-boundaries of interval modules, as well as their so-called facets, which are convenient characterizations of the interval supports.
Definition 12 (Upper- and lower-boundaries).
Given an interval $I\subset\mathbb{R}^{n}$, its upper-boundary $U[I]$ and lower-boundary $L[I]$ are defined as:
$L[I]:=\left\{x\in\bar{I}:\forall y\in\mathbb{R}^{n},\,y<_{n}x\Rightarrow y\not\in I\right\},\ \ U[I]:=\left\{x\in\bar{I}:\forall y\in\mathbb{R}^{n},\,y>_{n}x\Rightarrow y\not\in I\right\}.$
Moreover, the boundary of I{I} can be decomposed with $\partial{I}=L[I]\cup U[I]$. See Figure 3 for an illustration.
Figure 3: Lower- and upper-boundaries of an interval in $\mathbb{R}^{2}$ (Definition 12); and birthpoints and deathpoints $b_{x}^{I}$ and $d_{x}^{I}$ (Definition 15) of a point $x\in\mathbb{R}^{2}$.
When interval modules are discretely presented (see Definition 8), their lower- and upper-boundaries are made of flat parts, which are the faces of the corresponding rectangles forming the interval. Hence, we call facets the subsets of the lower- and upper-boundaries that are included in some hyperplanes of $\mathbb{R}^{n}$.
Definition 13 (Facet).
A lower (resp. upper) facet of an interval $I\subset\mathbb{R}^{n}$ is an (n1)(n-1)-submanifold of $\partial{\rm supp}(I)$ written as $\left\{x\in\mathbb{R}^{n}\,:\,x_{i}=c\right\}\cap L[I]$ (resp. $\left\{x\in\mathbb{R}^{n}\,:\,x_{i}=c\right\}\cap U[I]$) for some $c\in\mathbb{R}$ and some dimension $i\in\llbracket 1,n\rrbracket$ that is called the facet codirection. In particular, the upper- and lower-boundaries of a discretely presented interval module is a locally finite union of facets.
Fibered barcode.
The fibered barcode (23, 47) is a centerpiece of our MMA algorithm, and is defined, given a p.f.d. multi-parameter persistence module $\mathbb{M}$, as a map that takes as input a line (or segment) $l$ in $\mathbb{R}^{n}$, and outputs the persistence barcode associated to the single-parameter persistence module obtained by restricting $\mathbb{M}$ along $l$. We formalize these concepts in the next definition.
Definition 14 ( Diagonal fibered barcode).
Let $\mathbb{M}$ be a p.f.d. multi-parameter persistence module. Given a set $L$ of diagonal lines in $\mathbb{R}^{n}$ (i.e., lines with direction vector $1,\dots,1]\in\mathbb{R}^{n}$), we let the $L$-fibered barcode (or fibered barcode for short when $L$ is clear) be the family of barcodes associated to restrictions of the module along lines in $L$, i.e., ${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}={\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\right)_{l\in L}}$.
Remark 2.
Recall from (24, Theorem 2.8) that ${\rm supp}({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}})$ is a multiset of bars called persistence barcode: ${\rm supp}({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}})={\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)=\{[b_{i},d_{i})\}_{i\in\mathcal{I}(l)}$, where $\mathcal{I}(l)$ is an index set that depends on $\mathbb{M}$ and $l$. Moreover, when $\mathbb{M}=\bigoplus_{i\in\mathcal{I}}k^{I_{i}}$ is decomposable into interval modules, there are as many bars in the barcode as there are interval summands intersecting the line $l$: $|{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)|=|\{I_{i}:I_{i}\cap l\neq\varnothing\}|$.
It is also useful to characterize the fibered barcode with endpoints of lines.
Definition 15 (Birthpoint, Deathpoint).
Given a positive line $l$ (that is, a line whose direction vector $u$ is in $\mathbb{R}^{n}_{+}\setminus\left\{0\right\}$) and an interval $I\subseteq\mathbb{R}^{n}$, the birthpoint (resp. deathpoint) of $I$ along $l$ is:
$b^{I}_{l}:=\inf\,l\cap I,\quad\text{resp. }d^{I}_{l}:=\sup\,l\cap I.$
If the direction $u\in\mathbb{R}_{+}^{n}$ of the line is given by the context, and $x\in\mathbb{R}^{n}$, we will also let $b^{I}_{x}:=b_{l_{x}}^{I}$ (resp. $d^{I}_{x}:=d_{l_{x}}^{I}$) denote the birthpoint (resp. deathpoint) associated to $I$ and $x$, where $l_{x}$ is the line crossing $x$, with direction vector $u$. See Figure 3.
Remark 3 (Slicing an interval decomposable module).
Using birthpoints and deathpoints, the $L$-fibered barcode of an interval decomposable multi-parameter persistence module $\mathbb{M}=\bigoplus_{i\in\mathcal{I}}k^{I_{i}}$ can be written as:
${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}=\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\right)_{l\in L}=\left(\{[b^{I_{i}}_{l},d^{I_{i}}_{l})\}_{i\in\mathcal{I}}\right)_{l\in L}.$ (6)
Remark 4.
Notice that in Remark 3, the persistence barcodes ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)=\{[b^{I_{i}}_{l},d^{I_{i}}_{l})\}_{i\in\mathcal{I}}$ can be seen as multisets of segments $b_{i},d_{i})$ in $\mathbb{R}^{n}\cup\left\{\boldsymbol{\infty}{}\right\}$. In particular, the diagonal line of a given segment $b_{i},d_{i})$ can be recovered from, for instance, the birthpoint $b_{i}$, and hence, without loosing any information (except for lines with trivial barcodes), we will consider the following identification:
${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}=\bigcup_{l\in L}{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)=\left\{{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right):l\in L\right\}.$ (7)
Geometry and stability of diagonal barcodes.
We now present two simple yet fundamental results on diagonal barcodes. The first one characterizes rectangles formed by endpoints.
Lemma 1.
Let $l_{1},l_{2}$ be two diagonal lines and $k^{I}$} be a f.p. interval module such that the barcodes ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)$ are not empty. Let ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)=[b_{1},d_{1})$ and ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)=[b_{2},d_{2})$. Then, the rectangles $R_{b_{1},b_{2}}$, $R_{b_{2},b_{1}}$, $R_{d_{1},d_{2}}$ and $R_{d_{2},d_{1}}$ are flat, that is, they either have null volume, or their corners are not comparable.
Proof.
This lemma is a simple consequence of the persistence module definition: if $b_{1}$ and $b_{2}$ were comparable (as in Figure 4), then the rectangle $R_{b_{1},b_{2}}$ would not be trivial, and $b_{2}$ would not be a birthpoint since it would be possible to find a smaller birthpoint $\tilde{b}_{2}\leq_{n}b_{2}$ w.r.t. the partial order of $\mathbb{R}^{n}$ along the diagonal line passing through $b_{2}$. A similar argument holds for $d_{1}$ and $d_{2}$. See Figure 4.
Figure 4: Two bars $b_{1},d_{1})$ and $b_{2},d_{2})$ of an interval module.
We now show that endpoints of bars in barcodes associated to lines that are close should also be close. In other words, bars of the fibered barcode that are associated to lines that are close to each other must have similar length, as stated in the lemma below; see also (43, Lemma 2).
Lemma 2.
Let $k^{I}$} be a f.p. interval module, let $l_{1},l_{2}\subseteq\mathbb{R}^{n}$ be two diagonal lines and let $\overrightarrow{u}\in\mathbb{R}^{n}$ be a positive or negative vector such that $l_{2}=l_{1}+\overrightarrow{u}$. Then, the following properties hold:
  1. (i)
    If the barcode ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)=\{[b^{I}_{l_{1}},d^{I}_{l_{1}})\}$ is not empty and satisfies $\left\lVert d^{I}_{l_{1}}-b^{I}_{l_{1}}\right\rVert_{\infty}>\left\lVert\overrightarrow{u}\right\rVert_{\infty}$, then the barcode ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)$ is not empty as well, and
  2. (ii)
    If the barcodes ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)$ are not empty, then one has
    $\left\lVert d^{I}_{l_{1}}-d^{I}_{l_{2}}\right\rVert_{\infty}\leq\left\lVert\overrightarrow{u}\right\rVert_{\infty}\text{ and }\left\lVert b^{I}_{l_{1}}-b^{I}_{l_{2}}\right\rVert_{\infty}\leq\left\lVert\overrightarrow{u}\right\rVert_{\infty},$
    where we used the conventions $(+\boldsymbol{\infty})-(+\boldsymbol{\infty})=(-\boldsymbol{\infty})-(-\boldsymbol{\infty})=0$.
Proof.
Item $(i)$. Since $|(d^{I}_{l_{1}})_{i}-(b^{I}_{l_{1}})_{i}|=\left\lVert d^{I}_{l_{1}}-b^{I}_{l_{1}}\right\rVert_{\infty}>\left\lVert\overrightarrow{u}\right\rVert_{\infty}$ for any index $i\in\llbracket 1,n\rrbracket$, it follows that $b^{I}_{l_{1}}\leq_{n}b^{I}_{l_{1}}+\overrightarrow{u}\leq_{n}d^{I}_{l_{1}}$. [^4] We assume here that $\overrightarrow{u}$ is positive. It should be replaced by $d_{l_{1}}^{I}-\overrightarrow{u}$ if it is negative. Thus $b^{I}_{l_{1}}+\overrightarrow{u}$ must belong to I{I} since $I$ is an interval. Hence, since $b^{I}_{l_{1}}+\overrightarrow{u}\in l_{2}$, one has ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)={I}\cap l_{2}\neq\varnothing$.
Item $(ii)$. If one of the endpoints is infinite, the result holds trivially as the other endpoint has to be infinite too, so we now assume that the endpoints of the bars are all finite. Without loss of generality, assume that $l_{2}=l_{1}+\overrightarrow{u}$ where $\overrightarrow{u}$ is positive. Now, since both $d_{l_{2}}^{I}$ and $d_{l_{1}}^{I}+\overrightarrow{u}$ belong to $l_{2}$, they are comparable, so one has either $d_{l_{2}}^{I}>_{n}d_{l_{1}}^{I}+\overrightarrow{u}$ or $d_{l_{2}}^{I}\leq_{n}d_{l_{1}}^{I}+\overrightarrow{u}$. However, the first possibility would lead to $d_{l_{2}}^{I}>_{n}d_{l_{1}}^{I}+\overrightarrow{u}>_{n}d_{l_{1}}^{I}$, hence $d_{l_{1}}^{I}$ and $d_{l_{2}}^{I}$ would be (strictly) comparable in $\mathbb{R}^{n}$, which contradicts Lemma 1. Thus, one must have $d_{l_{2}}^{I}\leq_{n}d_{l_{1}}^{I}+\overrightarrow{u}$. Furthermore, and using the exact same arguments, $d^{I}_{l_{2}}-\overrightarrow{u}+\left\lVert\overrightarrow{u}\right\rVert_{\infty}\cdot\boldsymbol{1}$ is on $l_{1}$, and one must have $d_{l_{2}}^{I}-\overrightarrow{u}+\left\lVert\overrightarrow{u}\right\rVert_{\infty}\cdot\boldsymbol{1}\geq_{n}d_{l_{1}}^{I}$. Finally, by combining the two previous inequalities, one has:
$d_{l_{1}}^{I}-\left\lVert\overrightarrow{u}\right\rVert_{\infty}\cdot\textbf{1}\leq_{n}d_{l_{1}}^{I}+\overrightarrow{u}-\left\lVert\overrightarrow{u}\right\rVert_{\infty}\cdot\textbf{1}\leq_{n}d_{l_{2}}^{I}\leq_{n}d_{l_{1}}^{I}+\overrightarrow{u}\leq_{n}d_{l_{1}}^{I}+\left\lVert\overrightarrow{u}\right\rVert_{\infty}\cdot\textbf{1},$
which leads to the result for deathpoints. The proof extends straightforwardly to birthpoints. ∎
3 Computing candidate decompositions with the MMA algorithm
In this section, we present our family of descriptors for multi-parameter persistence modules, defined as candidate decompositions into interval summands and we identify a specific subfamily that we call approximate decompositions (Definition 19), in Section 3.1. Then, we show how practical computations of instances of such candidate decompositions can be done with our MMA algorithm in Sections 3.2 and 3.3.
3.1 Candidate and approximate decompositions
Our candidate decompositions depend on $\delta$-grids of lines, that we now define.
Definition 16 ($\delta$-grid of lines).
Let $K\subset\mathbb{R}^{n}$ be a compact set and $\delta>0$. The $\delta$-grid of lines associated to $K$, denoted as $L_{\delta}(K)$, is a family of diagonal lines evenly sampled in $K$:
$L_{\delta}(K):=\{l_{\delta\cdot u}:u\in\mathbb{Z}^{n}\text{ and }l_{\delta\cdot u}\cap K\neq\varnothing\},$
where $l_{\delta\cdot u}:=\delta\cdot u+\boldsymbol{e}_{\Delta}{\mathbb{R}}$ is the diagonal line with direction vector $\boldsymbol{e}_{\Delta}=[1,\dots,1]^{T}\in\mathbb{R}^{n}$ passing through $\delta\cdot u$.
Several new definitions can be introduced from grids of lines, which will turn useful either in the definition of our MMA algorithm, or in the corresponding theoretical proofs.
Definition 17 ($\delta$-regularly distributed lines filling a compact set).
Let $L$ be a set of diagonal lines in $\mathbb{R}^{n}$ and $K\subseteq\mathbb{R}^{n}$ be a compact set. Then, we say that:
  1. 1.
[…]
  2. 2.
    Two diagonal lines $l,l^{\prime}\in L$ are $\delta$-comparable if there exists a positive or negative vector $\overrightarrow{u}\in\mathbb{R}^{n}$ with $\left\lVert\overrightarrow{u}\right\rVert_{\infty}\leq\delta$ such that $l^{\prime}=l+\overrightarrow{u}$. If $\overrightarrow{u}$ is positive (resp. negative), we write $l^{\prime}\geq l$ (resp. $l^{\prime}\leq l$).
  3. 3.
    $L$ is $\delta$-regularly distributed if, for any pair of lines $(l,l^{\prime})\in L$, there exists a sequence of $\delta$-consecutive lines $\{l_{1},\dots,l_{k}\}$ in $L$ such that $l=l_{1}$ and $l^{\prime}=l_{k}$.
  4. 4.
    For a given line $l$ in a $\delta$-regularly distributed family of lines $L$, we call $L_{l}:=L\cap\{l+\delta\cdot\overrightarrow{u}:\overrightarrow{u}\in\{0,1\}^{n-1}\times\{0\}\}$ the $L$-surrounding set of $l$. In particular, one has $|L_{l}|\leq 2^{n-1}$.
  5. 5.
    $L$ $\delta$-fills $K$ (or fills $K$ when $\delta$ is clear) if any point of $K$ is at distance at most $\delta/2$ from some line in $L$. In other words, $K$ is included in the offset $L^{\delta/2}$.
Our candidate decompositions of a given multi-parameter persistence module $\mathbb{M}$ are, roughly speaking, interval decomposable modules with fibered barcodes containing the one of $\mathbb{M}$ on a $\delta$-grid of lines. Before going into the definition of our estimator candidate, we introduce a compactness assumption, that directly follows Remark 1.
Definition 18 (Module compactness).
We say that a rectangle $K=[a_{1},b_{1}]\times\cdots\times[a_{n},b_{n}]\subseteq\mathbb{R}^{n}$} compactly characterizes an $n$-parameter persistence module $\mathbb{M}$ if restricting $\mathbb{M}$ to $K$ preserve information, or, more formally, if $\mathrm{Lan}_{n}\,{\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{K}}\simeq\mathbb{M}$.
Definition 19 (Candidate and approximate decompositions).
Let $\mathbb{M}$ be a f.p. $n$-parameter persistence module. Let $K$ be a compact set that compactly characterizes $\mathbb{M}$, and $L:=L_{\delta}(K^{\delta})$ be the $\delta$-grid of lines of the offset $K^{\delta}=\{x\in\mathbb{R}^{n}:d_{\infty}(x,K)\leq\delta\}$, where $d_{\infty}$ stands for the $\left\lVert\cdot\right\rVert_{\infty}$ distance. A multi-parameter persistence module $\tilde{\mathbb{M}}_{\delta}$ is called a $\delta$-candidate decomposition of $\mathbb{M}$ if:
  1. $(i)$
    $\tilde{\mathbb{M}}_{\delta}$ is interval decomposable: $\tilde{\mathbb{M}}_{\delta}=\bigoplus_{i\in{\tilde{\mathcal{I}}}}{k^{\tilde{I}_{i}}}$, and
  2. $(ii)$
    ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\subseteq{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}_{\delta}\vphantom{\big|}\right|_{l}}\right)$ for any $l\in L$, i.e., the $L$-fibered barcode of $\mathbb{M}$, seen as a multiset of segments in $\mathbb{R}^{n}$, is included in the one of $\tilde{\mathbb{M}}_{\delta}$.
Clearly, a candidate decomposition can be a rough descriptor of $\mathbb{M}$, as the bars in its fibered barcode can be arbitrarily large. Hence, we identify a more stable subfamily of candidate decompositions:
  1. $(iii)$
    If ${d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}_{\delta}\vphantom{\big|}\right|_{l}}\right)\right)\leq 2\delta$ for any diagonal line $l$ (not only those that belong to $L$), then $\tilde{\mathbb{M}}_{\delta}$ is called an approximate decomposition of $\mathbb{M}$.
The reason we focus on preserving the diagonal fibered barcodes (instead of controlling, e.g., the rank invariant or the interleaving distance to $\mathbb{M}$) is because of the impossibility for general multi-parameter persistence modules to build a decomposition into indicator modules that is consistent with the rank invariant (see (46, Section 10.2.3)). Note however that this is still stronger than preserving the Hilbert function, i.e., the pointwise dimension of the module.
Remark 5 (Non-diagonal lines).
Extending the definition of our candidate decompositions to grids of non-diagonal lines (i.e., with direction vector different than $\boldsymbol{e}_{\Delta}$) is straightforward, and is completely equivalent to rescaling the filtrations. Using such non-diagonal grids will however produce less stable descriptors, as the interleaving distance (Definition 4) is based on the diagonal direction.
Remark 6.
One can check that the $\delta$-grid of lines of associated to $K^{\delta}$ used in Definition 19 is $\delta$-regularly distributed and $\delta$-fills $K$.
Finally, we introduce the definition of matching functions. Such functions play a key role in our MMA algorithm for computing candidate decompositions.
Definition 20 (Matching function).
Let $\mathbb{M}$ be a f.p. $n$-parameter persistence module, and $l,l^{\prime}\subseteq\mathbb{R}^{n}$ be two positive lines. A map $\sigma$ between the persistence barcodes:
$\sigma\colon{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\rightarrow{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right)\cup\left\{\varnothing\right\}$
is called an ($\mathbb{M}$-)matching function between $l$ and $l^{\prime}$ if the restriction of $\sigma$ to $\sigma^{-1}({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right))$ is injective. In other words, $\sigma$ is a partial bijection (Definition 5) between the two barcodes, seen as multisets of intervals.
Definition 21 (Induced matching functions).
If $\mathbb{M}=\bigoplus_{i\in\mathcal{I}}k^{I_{i}}$ is a f.p. interval decomposable module, then, for any positive line $l$, the bars of any barcode ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\cong\bigoplus_{i\in\mathcal{I}}{\left.\kern-1.2ptk^{{I_{i}}}\vphantom{\big|}\right|_{l}}$ can be indexed using $\mathcal{I}$ (by also counting empty bars). Thus, given any two positive lines $l_{1},l_{2}$, one can match the bars $\left.\kern-1.2ptk^{I_{i}}\vphantom{\big|}\right|_{l_{1}}$} and $\left.\kern-1.2ptk^{{I_{i}}}\vphantom{\big|}\right|_{l_{2}}$} together so that matched bars correspond to the same underlying interval summand of $\mathbb{M}$. In that case, the corresponding matching function $\sigma_{\mathbb{M}}$ is referred to as induced from $\mathbb{M}$.
3.2 Motivation for the MMA algorithm
Our MMA algorithm can be roughly described as a method that constructs interval summands based on families of bars (coming from the fibered barcode) that have been matched together using some matching function. The goal of this section is to frame the general question of practically computing candidate decompositions of a multi-parameter persistence module from its fibered barcode and a matching function. There are many ways of doing so, but the most natural ones are not necessarily the easiest computable ones. For the sake of simplicity, let us leave the problem of finding proper matching functions aside for now (which we will discuss in more details in Section 6), and assume that the underlying module is a single interval module $\mathbb{M}=\mathbb{I}$. Since interval modules are characterized by their supports, the goal is to recover ${\rm supp}(\mathbb{I})$. Moreover, if $\mathbb{I}$ is discretely presented, only the facets and critical points (i.e., points where several facets intersect) of ${\rm supp}(\mathbb{I})$ need to be captured or approximated. There are many different ways, for a given interval module $\mathbb{I}$, to define candidate critical points, that we call corners, using the endpoints of its fibered barcode, e.g., by using the minimum and maximum of consecutive endpoint coordinates. Hence, it is natural to find a candidate decomposition (or candidate interval in this case, since there is just one interval summand) $\tilde{\mathbb{I}}$ with model selection, i.e., by minimizing some penalty cost $\mathrm{pen}\colon S\to\mathbb{R}_{+}$, where $S$ is the set of discretely presented interval modules having the same fibered barcode than $\mathbb{I}$, or a subset thereof. See Figure 5 for examples of sets $S$ and corresponding candidate intervals. This penalty would forbid, e.g., overly complicated intervals that have lots of corners. For instance, minimizing the penalty:
[…]
Figure 5: Example of candidate decompositions for a 22-interval module $\mathbb{I}$ with support in $\mathbb{R}^{2}$. (Left) Given the $L$-fibered barcode of $\mathbb{I}$, where $L$ is the family of the four black lines, we want to approximate $\mathbb{I}$ with an element of $S$, i.e., an interval module with the same fibered barcode. (Middle) When one further constrains the set $S$ by asking to have at most one corner between two consecutive endpoints of the fibered barcode, the whole set $S$ can be computed explicitly. (Right) The set $S$ can also be described as the set of intervals which have to go through the blue path, and which can arbitrarily choose between the red or green path at three different locations. Hence, the cardinality of $S$ is $2^{3}$.
For interval modules, $S$ is generally a set of cardinal $c^{d}$, where $c$ is the number of candidate corners between birthpoints or deathpoints, and $d$ is the number of corners. For instance, in Figure 5, one has $n=2$, $c=2$ and $d=3$. Unfortunately, $c$ is of the order of $2^{n-1}$, and thus grows exponentially with the dimension $n$, and $d$ is difficult to control in practice, since it heavily depends on the number of lines in the fibered barcode and the regularity of the underlying interval module $\mathbb{I}$. Minimizing a penalty over $S$ is thus practical only for low dimension $n$ and small number of lines in the fibered barcode. Hence, our algorithm MMA presented in Section 3.3 does not use penalty minimization, but is rather defined with natural and simple corner choices.
Remark 7.
Note also that there are cases when the corner choices are canonical. For instance, any $2$-persistence module $\mathbb{M}$ with transition maps $\varphi_{\cdot}^{\cdot}$ that are weakly exact, i.e., that satisfy, for any $x\leq y$:
$\mathrm{im}\left(\varphi_{x}^{y}\right)=\mathrm{im}\left(\varphi_{(y_{1},x_{2})}^{y}\right)\cap\mathrm{im}\left(\varphi_{(x_{1},y_{2})}^{y}\right)\text{ and }\ker\left(\varphi_{x}^{y}\right)=\ker\left(\varphi_{x}^{(y_{1},x_{2})}\right)+\ker\left(\varphi_{x}^{(x_{1},y_{2})}\right),$
is rectangle decomposable (10). Hence, a canonical approximation of a summand $\mathbb{I}$ of $\mathbb{M}$ is given by the interval module whose support is the rectangle with corners $(\min_{l}(b^{I}_{l})_{1},\min_{l}(b^{I}_{l})_{2})$ and $(\max_{l}(b^{I}_{l})_{1},\max_{l}(b^{I}_{l})_{2})$, where $l$ goes through the family of lines $L$ of the fibered barcode.
3.3 The MMA algorithm for computing candidate decompositions
In this section, we introduce MMA: a fast algorithm for computing $\delta$-candidate decompositions. The pseudo-code for MMA is provided in Algorithm 1. Roughly speaking, given a f.p. $n$-parameter persistence module $\mathbb{M}$, an approximation parameter $\delta>0$, a $\delta$-grid of lines $L=L_{\delta}(K^{\delta})$ where $K$ compactly characterizes $\mathbb{M}$, and a matching function $\sigma$ (see Section 6 for a discussion about how to find such matching functions), Algorithm 1 works in three steps:
  • Step 1:
    compute the $L$-fibered barcode of $\mathbb{M}$,
  • Step 2:
    match together bars using the matching function $\sigma$,
  • Step 3:
    for each summand, use the endpoints of the corresponding bars to compute estimates of the critical points, using Algorithm 2.
Step 1 can be performed using any persistent homology software (such as, e.g., Gudhi, Ripser, Phat, etc), or with Rivet (47) when $n=2$. Our code is part of the multipers library (50), and can be found at https://github.com/DavidLapous/multipers. Moreover, it uses the vineyard algorithm (28), which allows us to run Steps 1 and 2 jointly (see Section 6.2).
Input 1: A f.p. $n$-parameter persistence module $\mathbb{M}$} and a compact $K\subset\mathbb{R}^{n}$ that compactly characterizes $\mathbb{M}$,
Input 2: $\delta$-grid of evenly spaced diagonal lines $L=L_{\delta}(K^{\delta})$
Input 3: Matching function $\sigma$
Output: Candidate decomposition $\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}$
Compute ${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}$, i.e., the $L$-fibered barcode of $\mathbb{M}$;
$\leftarrow []; # SS$ is the set of interval summands of the output candidate decomposition, intialized as the empty set
for $l\in L$ do
 for $b_{l}^{\mathbb{M}},d_{l}^{\mathbb{M}}]\in{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)$ do
    # Check whether it is in the image of the input matching
    if $\exists B\in S$ and $b,d]\in B$ s.t. $b_{l}^{\mathbb{M}},d_{l}^{\mathbb{M}}]=\sigma([b,d])$ then
       $B$.append($b_{l}^{\mathbb{M}},d_{l}^{\mathbb{M}}]$); # If it is, attach the bar to the corresponding summand
    # Otherwise initialize a new summand with the bar
    else
         Add $B:=[[b_{l}^{\mathbb{M}},d_{l}^{\mathbb{M}}]]$ to $S$;
# For each summand in $S$ characterized by a set of bars, build an approximate interval summand
Return $\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}:=\bigoplus_{B\in S}$ ApproximateInterval(B)(B);
Algorithm 1 MMA: Multi-parameter persistence Module Approximation.
We now describe the algorithm ApproximateInterval, which is used at the end of Algorithm 1. Its pseudo-code is given in Algorithm 2, and is defined in two steps:
  1. 1.
    first, we label birthpoints and deathpoints to identify facets with LabelEndpoints (Algorithm 3),
  2. 2.
    then, we use these labels to compute candidate critical points with ComputeCorners (Algorithm 4).
Input: Set of bars $B=\{[b_{l},d_{l})\}_{l\in L_{B}}$, where $L_{B}\subseteq L$
Output: Discretely presented interval module $k^{\tilde{I}(B)}$}
labs $\leftarrow$ LabelEndpoints(B)(B);
$C_{B}^{L}(B),C_{D}^{L}(B)\leftarrow$ ComputeCorners(B,(B, labs));
$\tilde{I}(B)$ $\leftarrow$ $\bigcup_{c\in C_{B}^{L}(B)}\bigcup_{c^{\prime}\in C_{D}^{L}(B)}R_{c,c^{\prime}}$;
Return $k^{\tilde{I}(B)}$;
Algorithm 2 ApproximateInterval
We first describe LabelEndpoints. The core idea of this algorithm, whose pseudo-code is given in Algorithm 3, is, for a given bar in $I$ associated to a line $l\in L$, to look at the corresponding surrounding set $L_{l}$ (see item (4) in Definition 17). If there exists a hyperplane $H$ such that all endpoints in this surrounding set belong to $H$, we identify $H$ as a facet, and we label the bar with the codirection of $H$.
Input: Set of bars $B=\{[b_{l},d_{l})\}_{l\in L_{B}}$, where $L_{B}\subseteq L$
Output: List labs of labels for each endpoint in $B$
[…]
for $l\in L_{B}$ do
 if $\exists i\in\llbracket 1,n\rrbracket$ and $c_{i}\in\mathbb{R},$ such that $\forall l^{\prime}\in L_{l},\,(b_{l^{\prime}})_{i}=c_{i}$ then
[…]
Return labs;
Algorithm 3 LabelEndpoints
Note that endpoints can have zero or more than one label. For instance, an endpoint that belongs to the intersection of several facets might have multiple labels. However, if several labels are identified, they must be associated to different dimensions. See Figure 6 for examples of label assignments when the underlying interval module has rectangle support.
Figure 6: Example of birthpoint labelling for an interval module $I$ with rectangle support with three surrounding sets of lines $L_{l_{1}}$, $L_{l_{2}}$, $L_{l_{3}}$ associated to three lines $l_{1},l_{2},l_{3}$. The labels of $l_{1},l_{2},l_{3}$ that are identified correspond to the red, blue and grey colored facets of II respectively.
Finally, we describe ComputeCorners. The core idea of the algorithm, whose pseudo-code is given in Algorithm 4, is to use the labels identified by LabelEndpoints to compute corners, or critical point estimates, in the following way: if all birthpoints (resp. deathpoints) in a surrounding set have at least one associated facet, i.e., have a non-empty list of labels, then a candidate corner can be defined using the minimum (resp. maximum) of all birthpoints (resp. deathpoints) coordinates. We only present the pseudo-code for birthpoints since the code for deathpoints is symmetric and can be obtained by replacing minimum by maximum and -$\infty$ by $\infty$. Note that the correctness of MMA follows directly from how these corners are computed: it is clear from Algorithm 4 that, for any bar in the fibered barcode, the algorithm will produce a corner that is lower (resp. larger) w.r.t. the partial order $\leq_{n}$ than the birthpoint (resp. deathpoint) of the bar (excluding the trivial case of corners with infinite coordinates).
Input 1: Set of bars $B=\{[b_{l},d_{l}]\}_{l\in L_{B}}$, where $L_{B}\subseteq L$
Input 2: List labs of labels for each endpoint in $B$
Output: List of birth corners $C_{B}$
$C_{B}\leftarrow[]$;
for $l\in L_{B}$ do
 $B_{L_{l}}\leftarrow\{b_{l^{\prime}}:l^{\prime}\in L_{l}\cap L_{B}\}$; # Note that $B_{L_{l}}\subseteq B$ by construction
 # Check whether all birthpoints in the surrounding set belong to $K$
 if $B_{L_{l}}\subseteq K$ then
    # Compute birth corner if all the birthpoints are labelled
[…]
       $\{(j,c_{j})\,:\,j\in\mathcal{J}\}\leftarrow\bigcup_{b\in B_{L_{l}}}$ labs(b)(b); # $\mathcal{J}\subseteq\llbracket 1,n\rrbracket$ is the set of codirections
         Define $C^{l}\in\mathbb{R}^{n}$ as
  • $(C^{l})_{j}=c_{j}$ if $j\in\mathcal{J}$
  • $(C^{l})_{j}=\min{\left\{(b_{l^{\prime}})_{j}\,:\,l^{\prime}\in L_{l}\cap L_{B}\right\}}$ otherwise
[…]
     # If the birthpoints are not all labeled, keep the birthpoints themselves as corners
    else
       for $l^{\prime}\in L_{l}\cap L_{B}$ do
[…]
 # If some birthpoints are not in $K$, they must correspond to infinite facets
 else
    Assert $B_{L_{l}}\cap K^{\delta}\backslash K\neq\varnothing$;
[…]
    $\{(j,c_{j})\,:\,j\in\mathcal{J}\}\leftarrow\bigcup_{b\in B_{L_{l}}}\textsc{labs}(b)$; # The cardinality of $\mathcal{J}$ must be strictly less than $n$
      Define $C^{l}\in\mathbb{R}^{n}$ as:
  • $(C^{l})_{j}=c_{j}$ if $j\in\mathcal{J}$
  • $(C^{l})_{j}=$-$\infty$ otherwise
[…]
Return $C_{B}$;
Algorithm 4 ComputeCorners
Complexity.
Computing the $L$-fibered barcode ${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}$ on a simplicial complex, as well as assigning the corresponding bars to their associated summands in the decomposition of $\mathbb{M}$, can be done with the vineyard algorithm (28) as matching function with complexity $O(N^{3}+|L|\cdot N\cdot T)$, where $N$ is the number of simplices in the simplicial complex, and $T$ is the maximal number of transpositions required to update the single-parameter filtrations corresponding to the consecutive lines in $L$. In the worst-case scenario, one has $T=N^{2}$. Note that $T$ usually decreases to a fixed constant as $|L|$ increases, and that this computation can be easily parallelized in practice.
Now, adding the complexities of Algorithms 3 and 4, the final complexity of Algorithm 1 is:
$O(N^{3}+|L|\cdot N\cdot T+|L|\cdot n\cdot 2^{n-1}).$ (9)
Of importance, the dependence on $n$ is much better than the (exact) decomposition algorithm proposed in (34, 32) whose complexity is $O(N^{n(2\omega+1)})$, where $\omega<2.373$ is the matrix multiplication exponent. It is also comparable to Rivet (47) (although Rivet only works when $n=2$), whose complexity is $O(N^{3}\kappa+(N+{\rm log}\kappa)\kappa^{2})$, where $\kappa=\kappa_{x}\kappa_{y}$ is the product of $x$ and $y$ coordinates used to evaluate the module (note that $\kappa_{x},\kappa_{y}$ are also user-dependent). The elder-rule staircode (17) works only for point cloud data when $n=2$ and homology dimension 0, but has better complexity $O(m^{2}\log(m))$, where $m$ is the number of points. Finally, note that our complexity can be controlled by the number of lines, which is user-dependent. We illustrate this useful property in Section 7.
Remark 8.
For the sake of simplicity and efficiency, the code that we provide at https://github.com/DavidLapous/multipers contains a simpler version of Algorithm 4, that does not compute and use labels, but rather gathers the birthpoints and deathpoints as corners directly. One can easily check that our approximation guarantees (Proposition 3 and Proposition 5) carry over to that simpler algorithm, however the exactness result (Proposition 12) is only valid for corners computed with Algorithm 4.
4 Theoretical robustness of MMA
In this section, our goal is to prove our first important result, Proposition 3, which states that, if the matching function is compatible, then the candidate decompositions computed by MMA are also approximate decompositions: they preserve the diagonal barcodes (up to $2\delta$) associated to all diagonal lines. We provide a proof in Section 4.1. We also discuss the stability of MMA w.r.t. $d_{\rm I}$} in Section 4.2.
4.1 Approximation guarantee of MMA
We first introduce so-called compatible matching functions, which are key elements for proving the approximation property satisfied by our MMA algorithm.
Definition 22 (Compatible matching function).
Let $\mathbb{M}$ be a f.p. $n$-parameter persistence module, and let $l_{1},l_{2}\subseteq\mathbb{R}^{n}$ be two diagonal lines that are at distance $\delta$ from each other. Assume ${\rm supp}(\mathbb{M})\cap l_{1}$ and ${\rm supp}(\mathbb{M})\cap l_{2}$ are not empty, and let $b_{1},d_{1})$ and $b_{2},d_{2})$ be bars in ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{1}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{2}}}\right)$, characterized by their endpoints. These bars are compatible if the rectangles $R_{b_{1},b_{2}}$, $R_{b_{2},b_{1}}$, $R_{d_{1},d_{2}}$ and $R_{d_{2},d_{1}}$ are flat or empty. Equivalently, two bars are compatible if their birthpoints (resp. deathpoints) are not strictly comparable, i.e., $b_{1}\not<_{n}b_{2}$, and $b_{1}\not>_{n}b_{2}$ (resp. $d_{1}\not<_{n}d_{2}$, and $d_{1}\not>_{n}d_{2}$). Moreover, we say that $b^{\mathbb{M}}_{l_{1}},d^{\mathbb{M}}_{l_{1}})$ is compatible with the empty set in $l_{2}$ if $\left\lVert b^{\mathbb{M}}_{l_{1}}-d^{\mathbb{M}}_{l_{1}}\right\rVert_{\infty}\leq 2\delta$.
A compatible matching function is a matching function that only pairs bars that are compatible.
Remark 9.
Induced matching functions (see Definition 21) are compatible, as per Lemma 1.
Proposition 3 (Approximation result).
Let $\mathbb{M}$ be a f.p. $n$-parameter persistence module and $\delta>0$. Let $K$ be a rectangle in $\mathbb{R}^{n}$ that compactly characterizes $\mathbb{M}$, and $L:=L_{\delta}(K^{\delta})$ be the $\delta$-grid of lines of the offset $K^{\delta}$. Finally, let $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}:=\mathtt{MMA}{}({\mathbb{M}},L,\sigma)$}, where $\sigma$ is a compatible matching function. Then $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}$ is a $\delta$-approximate decomposition of $\mathbb{M}$. More precisely, given some diagonal line $l$, one has:
  1. (i{i})
    ${d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\emph{{MMA}{}}}_{\delta}\vphantom{\big|}\right|_{l}}\right)\right)=0$ if $l\in L$, and
  2. (ii{ii})
    ${d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\emph{{MMA}{}}}_{\delta}\vphantom{\big|}\right|_{l}}\right)\right)\leq 2\delta$ otherwise.
In order for Proposition 3 to apply, one needs to find a compatible matching function $\sigma$. We discuss how to design such matching functions for interval decomposable modules in Section 6.1 and for general $2$-parameter modules in Section 6.2. We also hypothesize that compatible matching functions for general $n$-parameter persistence modules could be constructed using representative cycles in a similar way than the construction of the graphcode (39, 40), a conjecture that we leave for future work.
Proof.
We first prove (i{i}). Let $l\in L$, and let $b_{l}$ be the birthpoint of a bar $b$ in ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)$. Let $B$ be the set of bars containing $b$ computed with Algorithm 1, let $C_{B}^{L}(B)$ and $C_{D}^{L}(B)$ be the birth and death corners computed with Algorithm 4, and let $\tilde{I}$ be the interval computed with Algorithm 2, i.e., one has:
$\tilde{I}=\bigcup_{c\in C_{B}^{L}(B)}\bigcup_{c^{\prime}\in C_{D}^{L}(B)}R_{c,c^{\prime}},$ (10)
In order to show (i{i}), we first need to show that $b_{l}=b^{\tilde{I}}_{l}$ (and then the proof for deathpoints will follow by symmetry), where $b^{\tilde{I}}_{l}$ is defined as per Definition 15. Note that $b_{l}$ and $b^{\tilde{I}}_{l}$ are comparable since they belong to the same diagonal line $l$.
Strategy. In order to show $b_{l}=b^{\tilde{I}}_{l}$, we are going to show that 1. $b_{l}\leq_{n}b^{\tilde{I}}_{l}$ and 2. $b^{\tilde{I}}_{l}\leq_{n}b_{l}$.
  1. 1.
    In order to show $b_{l}\leq_{n}b^{\tilde{I}}_{l}$, we are going to show that $c\not<_{n}b_{l}$ for any corner $c\in C_{B}^{L}(B)$. Indeed, if one assumes $b_{l}>_{n}b^{\tilde{I}}_{l}$ by contradiction, and since there always exists a birth corner $c\in C_{B}^{L}(B)$ such that $c\leq_{n}b^{\tilde{I}}_{l}$ by construction of $\tilde{I}$, one has $c\leq_{n}b^{\tilde{I}}_{l}<_{n}b_{l}$.
  2. 2.
    In order to show $b^{\tilde{I}}_{l}\leq_{n}b_{l}$, we are going to show that there exists a corner $c\in C_{B}^{L}(B)$ such that $c\leq_{n}b_{l}$. Indeed, if there is such a birth corner, and if $b^{\tilde{I}}_{l}>_{n}b_{l}$ by contradiction, then $c\leq_{n}b_{l}<_{n}b^{\tilde{I}}_{l}$, and $R_{c,b^{\tilde{I}}_{l}}$ is not flat, contradicting Lemma 1.
Proof of (2). By construction of $\tilde{I}$ with Algorithm 4, if $b_{l}$ is labelled, then there exists a line $l^{\prime}$ and a corner $c^{l^{\prime}}\in C_{B}^{L}(B)$ that is smaller than $b_{l}$ so we can take $c:=c^{l^{\prime}}$. If $b_{l}$ is not labelled, it belongs itself to $C_{B}^{L}(B)$, and we can take $c:=b_{l}$.
Proof of (1). Let $c\in C_{B}^{L}(B)$ be a birth corner, and let $L_{l_{0}}$ be the associated surrounding set of lines for some $l_{0}\in L$. Let $c]_{l}:=\min\left[\left(c+\left(\mathbb{R}_{+}\right)^{n}\right)\cap l\right]$ be the smallest element in the intersection between the positive cone on $c$ and $l$. Assume $c]_{l}\geq_{n}b_{l}$ and $c<_{n}b_{l}$. Then $R_{c,[c]_{l}}$ is not flat, contradicting the fact that $c]_{l}$ is the smallest element. Thus, we only have to show $c]_{l}\geq_{n}b_{l}$. There are two cases.
  1. 1.
    Either some birthpoints of $L_{l_{0}}$ are not labelled by Algorithm 3, and $c$ is equal to the birthpoint $b_{l^{\prime}}$ of another bar in ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right){\cap B}$ for some $l^{\prime}\in L_{l_{0}}$. Now, assume $c]_{l}<_{n}b_{l}$ by contradiction. Then $b_{l^{\prime}}=c\leq_{n}[c]_{l}<_{n}b_{l}$. Thus $b_{l^{\prime}}<_{n}b_{l}$ and $R_{b_{l^{\prime}},b_{l}}$ is not flat, contradicting the fact that $\sigma$ is compatible. Hence $c]_{l}\geq_{n}b_{l}$.
  2. 2.
    Or all the birthpoints of $L_{l_{0}}$ are labelled by Algorithm 3. Again, we study two separate cases. See Figure 7 for an illustration.
    1. (a)
      Either $l\in L_{l_{0}}$. Then, $\exists i\in\llbracket 1,n\rrbracket$ such that $(b_{l})_{i}=c_{i}$. This yields $(b_{l})_{i}=c_{i}\leq([c]_{l})_{i}$, and thus $c]_{l}\geq_{n}b_{l}$ since they both belong to the same diagonal line $l$.
    2. (b)
      Or the line $l$ does not belong to $L_{l_{0}}$. Since $c]_{l}$ is on the boundary of the positive cone based on $c$, there exists $i\in\llbracket 1,n\rrbracket$ such that $([c]_{l})_{i}=c_{i}$. Assume again by contradiction that $b_{l}>_{n}[c]_{l}$, and write:
      $c]_{l}=c+\sum_{j\neq i}(\delta\alpha_{j})e_{j}=:c+\overrightarrow{v}<_{n}b_{l},$
      with $\alpha_{j}\geq 0$ for $j\in\llbracket 1,n\rrbracket\backslash\left\{i\right\}$. Since $l\notin L_{l_{0}}$, there exists some $j_{0}$ such that $\alpha_{j_{0}}>1$. Let $\overrightarrow{u}:=((\overrightarrow{v}_{j}\,\mathrm{mod}\,\delta)_{j\in\llbracket 1,n\rrbracket})=(([c]_{l}-c)_{j}\,\mathrm{mod}\,\delta)_{j\in\llbracket 1,n\rrbracket}\in[0,\delta)^{n}\leq_{n}\overrightarrow{v}$. Let $l^{\prime}:=l_{c+\overrightarrow{u}}$ be the diagonal line passing through $c+\overrightarrow{u}$. Now, recall that the lines of $L$ are drawn on a grid, so $l^{\prime}\in L$ since $l^{\prime}=l+\overrightarrow{u}-\overrightarrow{v}$. Moreover, one has: by definition, $c\in{\rm conv}(L_{l_{0}})$. Since the lines of $L$ are on a grid, one has:
      $\forall l_{1},l_{2}\in L,\quad\|l_{1}\cap H_{n},{\rm conv}(L_{l_{2}})\cap H_{n})\|_{\infty}<\delta\Longrightarrow l_{1}\in L_{l_{2}},$
      where $H_{n}=\left\{x\in\mathbb{R}^{n}:x_{n}=c_{n}\right\}$. Now, note that $c+\overrightarrow{u}$ and $c+\overrightarrow{u}-\overrightarrow{u}_{n}\cdot\textbf{1}$ both belong to $l^{\prime}$, and that $c+\overrightarrow{u}-\overrightarrow{u}_{n}\cdot\textbf{1}\in H_{n}$. Moreover, since:
      $\left\lVert(c+(\overrightarrow{u}-\overrightarrow{u}_{n}\cdot\textbf{1}))-c\right\rVert_{\infty}=\left\lVert\overrightarrow{u}-\overrightarrow{u}_{n}\cdot\textbf{1}\right\rVert_{\infty}<\delta,$
      one has $l^{\prime}\in L_{l_{0}}$. Thus, letting $b_{l^{\prime}}$ be the birthpoint of the corresponding bar in ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right)\cap B$, there exists $i^{\prime}\in\llbracket 1,n\rrbracket$ such that $(b_{l^{\prime}})_{i^{\prime}}=c_{i^{\prime}}\leq(c+\overrightarrow{u})_{i^{\prime}}$ and thus $b_{l^{\prime}}\leq_{n}(c+\overrightarrow{u})$ since $b_{l^{\prime}}$ and $c+\overrightarrow{u}$ are comparable on the diagonal line $l^{\prime}$. Finally, $b_{l^{\prime}}\leq_{n}c+\overrightarrow{u}\leq_{n}c+\overrightarrow{v}<_{n}b_{l}$, and $R_{b_{l^{\prime}},b_{l}}$ is not flat, contradicting the fact that $\sigma$ is compatible. Hence, $b_{l}\leq_{n}[c]_{l}$.
    Figure 7: Illustration of $l,l^{\prime},c,[c]_{l},[c]_{l^{\prime}},\overrightarrow{u},\overrightarrow{v},b_{l},b_{l^{\prime}}$ when one assumes that $c]_{l}<_{n}b_{l}$.
The proof applies straightforwardly to deathpoints by symmetry.
Now, the proof of (ii{ii}) is then a simple consequence of Lemma 2. Indeed, given a line $l\not\in L$, there must be a line $l^{*}\in L$ such that $l^{*}=l+\overrightarrow{u}$ with $\left\lVert\overrightarrow{u}\right\rVert\leq\delta$ since $L$ fills $K$. Then, one has:
$\begin{array}[]{lcl}{d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big|}\right|_{l}}\right)\right)&\leq&{d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{*}}}\right)\right)+{d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{*}}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big|}\right|_{l^{*}}}\right)\right)+{d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big|}\right|_{l^{*}}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big|}\right|_{l}}\right)\right)\\ &\leq&\delta+0+\delta=2\delta,\end{array}$
by Lemma 2. ∎
Instability of MMA w.r.t. interleaving distance $d_{\rm I}$}.
While using compatible matching functions helps controlling the diagonal fibered barcodes, it is unfortunately not sufficient for bounding the interleaving distance: outputs of MMA can be very far in terms of $d_{\rm I}$} while the modules they are computed from are not. We provide two multi-parameter persistence modules in Figure 8 that illustrate such lack of stability. In this figure, the two $2$-parameter filtrations only differ on the middle edge (in blue) of the simplicial complex. When the appearance of this edge is delayed (as is the case for the multi-parameter filtration displayed on top), the bars in the barcodes corresponding to the lower and upper cycles of the simplicial complex get paired by the compatible matching function, and create together the large red summand. On the other hand, this does not happen for the other multi-parameter filtration: the bars corresponding to these cycles are never paired and form distinct interval summands with same size.
Figure 8: (Left) Two bi-filtrations whose corresponding multi-parameter persistence modules in homology dimension 11 are $\varepsilon$-interleaved. (Right) The two corresponding, significantly different interval decompositions obtained with MMA computed with a compatible matching. Intervals in these decompositions are displayed with red and yellow colors.
Figure 8 illustrates the price to pay for designing interpretable decompositions (i.e., such that each summand corresponds to a cycle of the simplicial complex) when several choices are possible: there are several different ways to assign cycles to summands in the non-interval decomposable module displayed on top of Figure 8—the one computed by MMA (shown in the figure) being one possibility. An important conjecture of this article, that is left for future work, is that stacking the candidate decompositions produced by our MMA algorithm for all possible compatible matching functions induces a complete topological invariant of the module.
4.2 Stability property of MMA
As it is not possible to control powerful distances such as $d_{\rm I}$}, we end this section by ensuring that MMA can still remain stable w.r.t. to the data itself by appropriately choosing the matching functions. Indeed, given two multi-parameter filtrations computed from the sublevel sets of functions $f,g$, Equation (1) ensures that the bottleneck distances between barcodes in the fibered barcodes of ${\mathbb{M}}(f)$ and ${\mathbb{M}}(g)$ are upper bounded by $\left\lVert f-g\right\rVert_{\infty}$. This in turns means that we can fix an (arbitrary) compatible matching function $\sigma_{f}$ (if it exists) for computing an approximate decomposition of $f$ with MMA, and define another one $\sigma_{g}$ that commutes with $\sigma_{f}$ and the optimal partial matching $\nu$ given by those bottleneck distances. Doing so leads to the following proposition.
Proposition 4 (Enforced stability).
Let $f,g:S\to\mathbb{R}^{n}$ be two continuous functions defined on a topological space $S$, and let $\mathbb{M}(f)$ and $\mathbb{M}(g)$ be the multi-parameter persistence modules associated to the homology groups of their sublevel sets. Let $K$ be a rectangle in $\mathbb{R}^{n}$ that compactly characterizes $\mathbb{M}(f)$ and $\mathbb{M}(g)$, and $L:=L_{\delta}(K^{\delta})$ be the $\delta$-grid of lines of the offset $K^{\delta}$. Finally, let $\sigma_{f}$ be an arbitrary compatible matching function. Then, there exists a matching function $\sigma_{g}$, such that the following diagram commutes:
[…]
In particular, assuming that $\sigma_{g}$ is also compatible, if we let:
$\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(f):=\mathtt{MMA}(\mathbb{M}(f),L,\sigma_{f})\quad\textnormal{and}\quad\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(g):=\mathtt{MMA}(\mathbb{M}(g),L,\sigma_{g}),$ (12)
we have the following stability inequality:
${d_{\rm I}}(\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(f),\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(g))\leq{d_{\rm B}}(\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(f),\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(g))\leq\left\lVert f-g\right\rVert_{\infty}+\delta.$ (13)
Proof.
Let $\sigma_{g}$ be the matching function induced by the following commutative diagram:
${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}(f)\vphantom{\big|}\right|_{l}}\right)$${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}(f)\vphantom{\big|}\right|_{l^{\prime}}}\right)$${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}(g)\vphantom{\big|}\right|_{l}}\right)$${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}(g)\vphantom{\big|}\right|_{l^{\prime}}}\right),$$\scriptstyle{\sigma_{f}}$$\scriptstyle{\nu_{l,{d_{\rm B}}}}$$\scriptstyle{\nu_{l^{\prime},{d_{\rm B}}}}$$\scriptstyle{\sigma_{g}}$ (14)
where $\nu_{l,{d_{\rm B}}}$ denotes the optimal partial matching induced by ${d_{\rm B}}({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}(f)\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}(g)\vphantom{\big|}\right|_{l}}\right))$.
Let $k^{I_{f}},k^{I_{g}}$ denote two interval summands of $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(f)$ and $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(g)$ that are paired by $\nu_{\cdot,{d_{\rm B}}}$. Then, controlling the bottleneck distance between these outputs of MMA simply amounts to controlling the Hausdorff distance between $I_{f}$} and $I_{g}$}. In order to control this distance, let $B_{f}$ and $B_{g}$ denote the bars that induced these interval summands as per Algorithm 2. Then, one has $B_{f}\subseteq{I_{f}}$ and $B_{g}\subseteq B_{f}^{\gamma}$, where $\gamma=\left\lVert f-g\right\rVert_{\infty}$. Thus, since ${\rm supp}(I_{g})\subseteq B_{g}^{\delta}$ (by construction), one has $I_{g}\subseteq B_{f}^{\gamma+\delta}\subseteq{\rm supp}(I_{f})^{\gamma+\delta}$. The result follows by symmetry of $f$ and $g$. ∎
Note that finding compatible matching functions can be weakened into a convex problem (as compatibility can be checked with a sequence of inequalities), thus inducing an open question: is it possible to define an optimization problem whose minimization would yield a matching function that is both compatible and stable with respect to the input data? The paragraph at the end of Section 4.1 shows in particular that adding another approximation term is necessary if one wants to avoid decomposition instabilities.
5 The case of interval decomposable modules
In this section, we refine Proposition 3 to interval decomposable modules. In particular, we show that, upon using induced matching functions (see Definition 21), the approximate decompositions computed by our MMA algorithm become stable w.r.t. the interleaving and bottleneck distances (Proposition 5) in Section 5.1, and that the input interval decomposable module can even be recovered exactly for a small yet positive $\delta$ (Proposition 12) in Section 5.2.
5.1 Stability w.r.t. interleaving and bottleneck distances
The goal of this section is to show the following result:
Proposition 5 (Stability result).
Let $\mathbb{M}$ be a f.p. interval decomposable $n$-parameter persistence module. Let $K$ be a rectangle in $\mathbb{R}^{n}$ that compactly characterizes $\mathbb{M}$, and $L:=L_{\delta}(K^{\delta})$ be the $\delta$-grid of lines of the offset $K^{\delta}$. Finally, let $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}:=\mathtt{MMA}{}({\mathbb{M}},L,\sigma)$}, where $\sigma$ is a matching function that commutes with the induced matching function $\sigma_{\mathbb{M}}$. More precisely, denoting $\mathbb{M}=\bigoplus_{i\in\mathcal{I}}k^{I_{i}}$ and $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}=\bigoplus_{i\in\tilde{\mathcal{I}}}k^{\tilde{I}_{i}}$, this means that there exists a bijection $\nu\colon\mathcal{I}_{L}\to\tilde{\mathcal{I}}$, where $\mathcal{I}_{L}=\{i\in\mathcal{I}:I_{i}\cap L\neq\varnothing\}$, such that, for any two lines $l,l^{\prime}\in L$, the following diagram commutes:
${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)$${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right)$${\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}\vphantom{\big|}\right|_{l}}\right)$${\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}\vphantom{\big|}\right|_{l^{\prime}}}\right),$$\scriptstyle{\sigma_{\mathbb{M}}}$$\scriptstyle{\nu_{l}}$$\scriptstyle{\nu_{l^{\prime}}}$$\scriptstyle{\sigma}$
where $\nu_{l}\colon{\left.\kern-1.2ptI_{i}\vphantom{\big|}\right|_{l}}\in{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\mapsto{\left.\kern-1.2pt\tilde{I}_{\nu(i)}\vphantom{\big|}\right|_{l}}\in{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}\vphantom{\big|}\right|_{l}}\right)$ (and similarly for $l^{\prime}$).
Then, one has:
${d_{\rm I}}({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\emph{{MMA}{}}}_{\delta}\vphantom{\big|}\right|_{K}},{\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big|}\right|_{K}})\leq{d_{\rm B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\emph{{MMA}{}}}_{\delta}\vphantom{\big|}\right|_{K}},{\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big|}\right|_{K}})\leq\delta.$
One might wonder how to construct matching functions that commute with $\sigma_{\mathbb{M}}$ in practice. It turns out that any compatible matching function, as well as the matching functions associated to the Wasserstein distances and the vineyards algorithm, all commute with the induced matching $\sigma_{\mathbb{M}}$ for small enough $\delta$ and under some generic assumptions, as we show in Section 6.1.
Note also that it is possible to generalize Proposition 5 to modules that are not restricted to $K$ by constraining the parts of the candidate decompositions that are outside of $K$ with Kan extensions, but we stick to our formulation for the sake of simplicity. We will now prove a few technical results and lemmas in Section 5.1.1, and we finally prove Proposition 5 in Section 5.1.2.
5.1.1 Additional lemmas
In this section, we prove a few preliminary results about endpoints of interval modules, that will turn out useful for proving Proposition 5.
Endpoint location.
The next definition and result show that endpoints of an interval module must be located in the vicinity of the other endpoints of the module that are close to it—more precisely, in their rectangle hull. This will be useful in the proof of Proposition 5; in particular, we will use this result to characterize the positions of endpoints of any given diagonal line $l$ solely from the endpoints of the lines of the grid $L$ that are close to $l$.
Definition 23.
Let $S\subseteq\mathbb{R}^{n}$. The rectangle hull of $S$, denoted by $\mathrm{recthull}[S]$, is defined as the smallest rectangle containing $S$:
$\mathrm{recthull}[S]:=\left\{x\in\mathbb{R}^{n}:\forall i\in\llbracket 1,n\rrbracket,\min_{s\in S}s_{i}\leq x_{i}\leq\max_{s\in S}s_{i}\right\}=R_{\wedge S,\vee S},$
where $(\wedge S)_{i}:=\min_{s\in S}s_{i}$ and $(\vee S)_{i}:=\max_{s\in S}s_{i}$.
Lemma 6 (Endpoints bound).
Let $k^{I}$ be a f.p. interval module. Let $K$ be a rectangle in $\mathbb{R}^{n}$ and $L:=L_{\delta}(K^{\delta})$ be the $\delta$-grid of lines of the offset $K^{\delta}$. Let $x\in K$, $l_{x}$ be the diagonal line passing through $x$, and let $L_{x,\delta}:=\{l\in L\,:\,d_{\infty}(x,l)\leq\delta\text{ and }l_{x},l\text{ are $\delta$-comparable}\}$, which is non-empty since $L$ $\delta$-fills $K$. Assume that $l_{x}\cap I\neq\varnothing$, and $d^{I}_{x}\in U[I]$ be the associated deathpoint, and assume that for any line $l$ in $L_{x,\delta}$, one has $I\cap l\neq\varnothing$, and let $D_{x,\delta}^{I}$ be the set of the associated deathpoints: $D_{x,\delta}^{I}=\{d_{l}^{I}\,:\,l\in L_{x,\delta}\}$. Then, $d^{I}_{x}$ belongs to the rectangle hull of a subset $\tilde{D}_{x,\delta}^{I}$ of $D_{x,\delta}^{I}$: one has $d^{I}_{x}\in\mathrm{recthull}[\tilde{D}_{x,\delta}^{I}]$ with $\tilde{D}_{x,\delta}^{I}\subseteq D_{x,\delta}^{I}$.
Similarly, if $b^{I}_{x}\in L[I]$ is a birthpoint, then $b^{I}_{x}\in\mathrm{recthull}[\tilde{B}_{x,\delta}^{I}]$, where $\tilde{B}_{x,\delta}^{I}$ is a subset of $B_{x,\delta}^{I}=\{b_{l}^{I}\,:\,l\in L_{x,\delta}\}$, i.e., the set of birthpoints associated to $L_{x,\delta}$.
In other words, the endpoints of an interval module always belong to the rectangle hull of the endpoints associated to neighbouring lines. See Figure 9 for an illustration.
Figure 9: Example of deathpoint bound in $\mathbb{R}^{3}$, with $d\in U[I]$, and $D_{x,\delta}^{I}=\left\{d_{1},d_{2},d_{3},d_{4}\right\}$. (Left) Rectangle hull of the deathpoints $D_{x,\delta}^{I}$. (Right) Upper-boundary $U[I]$.
Proof.
We first prove the result for deathpoints. Note that the result is trivially satisfied if $d^{I}_{x}$ and the deathpoints in $D^{I}_{x,\delta}$ are infinite, so we assume that they are finite in the following. To alleviate notations, we let $d:=d^{I}_{x}$. Let $j\in\llbracket 1,n\rrbracket$ be an arbitrary dimension. In order to prove the result, we will show that there exist two deathpoints $\underline{d}$ and $\overline{d}$ associated to consecutive lines of $L_{x,\delta}$ such that $\underline{d}_{j}\leq d_{j}\leq\overline{d}_{j}$.
Construction of $\underline{d},\overline{d}$. Let $H_{j}$ be the hyperplane $H_{j}=d+e_{j}^{\perp}$. Since $L$ $\delta$-fills $K$, there exists a diagonal line $l\in L$ such that $d_{\infty}(x,l)\leq\delta/2$. Moreover, since $l$ and $l_{x}$ (the line passing through $x$ and $d$) are both diagonal, one has $d_{\infty}(d,l)=d_{\infty}(x,l)\leq\delta/2$. Let $\pi_{l}(d)\in l$ be the projection of $d$ onto $l$ that achieves $d_{\infty}(d,l)$, and let $d^{j}:=l\cap H_{j}$. See Figure 10 for an illustration of these objects.
Figure 10: Illustration of $H_{j},d,l,d^{j}$.
[…]
By construction, one has $d^{-}\leq d\leq d^{+}\in H_{j}$. and
$\|d^{+}-d\|_{\infty},\|d^{-}-d\|_{\infty}\leq\delta.$ (15)
Since $l$ and the diagonal lines $\underline{l}$ and $\overline{l}$ passing through $d^{-}$ and $d^{+}$ respectively are $\delta$-consecutive, and since $x\in K$, the projections of $x$ onto $\underline{l}$ and $\overline{l}$ are in $K^{\delta}$, and thus $\underline{l},\overline{l}$ must belong to $L$, and thus to $L_{x,\delta}$, as by construction $l_{x}$ is $\delta$-comparable with the diagonal lines $\underline{l}$ and $\overline{l}$. Let $\overline{d}:=d_{\underline{l}}^{I}\in\underline{l}$ and $\underline{d}:=d_{\overline{l}}^{I}\in\overline{l}$ be their deathpoints (which exist by assumption).
Proof of inequalities. We now show that $\overline{d}_{j}\geq_{n}d_{j}\geq_{n}\underline{d}_{j}$. We start with the second inequality. Since $d^{+}$ and $\underline{d}$ are one the same diagonal line, they are comparable. Furthermore, if one had $d^{+}<_{n}\underline{d}$ by contradiction, then the induced rectangle $R_{d,\underline{d}}$ would not be flat since $d\leq_{n}d^{+}<_{n}\underline{d}$, which would contradict Lemma 1. As a consequence, $d^{+}\geq_{n}\underline{d}$. Taking the $j$-th coordinate yields $d_{j}=d^{+}_{j}\geq\underline{d}_{j}$. The first inequality holds using the same arguments.
This proof applies straightforwardly to birthpoints by symmetry. ∎
Using Lemma 2, one can generalize Lemma 6 above to the case where some lines in $L_{x,\delta}$ have an empty intersection with $I$, and then define a common location for all endpoints that belong to the convex hull of the same $L$-surrounding set, as we do in the following proposition.
Proposition 7.
Let $k^{I}$} be a f.p. interval module. Let $K$ be a rectangle in $\mathbb{R}^{n}$ and $L:=L_{\delta}(K^{\delta})$ be the $\delta$-grid of lines of the offset $K^{\delta}$. Let $l\in L$ such that $|L_{l}|=2^{n-1}$ and assume that ${\rm conv}(L_{l})\cap L[I]$ (resp. ${\rm conv}(L_{l})\cap U[I]$) is not empty. Then, there exists a set $B_{l}$ (resp. $D_{l}$) such that for any $x\in{\rm conv}(L_{l})\cap L[I]$ (resp. ${\rm conv}(L_{l})\cap U[I]$), one has either $\|b_{x}^{I}-d_{x}^{I}\|_{\infty}\leq\delta$, or $x\in B_{l}$ (resp. $D_{l}$), where $B_{l}$ (resp. $D_{l}$) is a rectangular set in $\mathbb{R}^{n}$ that can be constructed from the birthpoints $(b_{l^{\prime}}^{I})_{l^{\prime}\in L_{l}}$ (resp. deathpoints $(d_{l^{\prime}}^{I})_{l^{\prime}\in L_{l}}$). Moreover, one has that:
${\rm diam}(B_{l})={\rm sup}_{x,x^{\prime}\in B_{l}}\left\lVert x-x^{\prime}\right\rVert_{\infty}\leq\delta,$ (16)
and similarly for $D_{l}$.
Proof.
We first construct $B_{l}$ and $D_{l}$, and then we will show items (1) and (2).
Definition of $B_{l},D_{l}$. Let first assume that $x$ is in the interior of ${\rm conv}(L_{l})$, that we denote with ${\kern 0.0pt{\rm conv}(L_{l})}^{\mathrm{o}}$. Note that if there is a line $l_{0}$ that is $\delta$-comparable to $l_{x}$, and such that ${\mathcal{B}}\left({\left.\kern-1.2ptk^{I}\vphantom{\big|}\right|_{l_{0}}}\right)=\varnothing$, then by Lemma 2 $(i)$, one immediately has $\|b_{x}^{I}-d_{x}^{I}\|_{\infty}\leq\delta$. Hence, we now assume that the barcodes along any line that is $\delta$-comparable to $l_{x}$ is not empty, which means that the hypotheses of Lemma 6 are satisfied for $x$. Now, remark that since $L$ is a grid, if one is able to find a line $l^{\prime}$ in $L$ whose intersections with hyperplanes associated to the canonical axes of $\mathbb{R}^{n}$ are $\delta$-close to $x$, then, since $x$ is in the interior of an $L$-surrounding set $L_{l}$, $l^{\prime}$ must belong to that surrounding set $L_{l}$ as well. More formally, one has that, for any line $l^{\prime}\in L$:
$d_{\infty}(x,l^{\prime}\cap H_{i})\leq\delta\quad\Longrightarrow\quad l^{\prime}\in L_{l},\quad\text{ where }H_{i}=\left\{y\in\mathbb{R}^{n}:y_{i}=x_{i}\right\}.$
This ensures (from Equation 15) that the lines of $L$ associated to $\tilde{B}^{I}_{x,\delta}$ and $\tilde{D}^{I}_{x,\delta}$ are all included in $L_{l}$ for any $x\in{\kern 0.0pt{\rm conv}(L_{l})}^{\mathrm{o}}$, and thus that we can safely define:
$D_{l}:=\overline{\bigcup_{x\in{\kern 0.0pt{\rm conv}(L_{l})}^{\mathrm{o}}}\mathrm{recthull}[\tilde{D}_{x,\delta}^{I}]}\quad\text{and}\quad B_{l}:=\overline{\bigcup_{x\in{\kern 0.0pt{\rm conv}(L_{l})}^{\mathrm{o}}}\mathrm{recthull}[\tilde{B}_{x,\delta}^{I}]}.$
Note that $B_{l}$ and $D_{l}$ depend only on the endpoints of the lines in $L_{l}$ and that $d_{x}^{I}\in D_{l}$ and $b_{x}^{I}\in B_{l}$ for any $x\in{\kern 0.0pt{\rm conv}(L_{l})}^{\mathrm{o}}$ by Lemma 6. Furthermore, if $x$ is in the closure of ${\rm conv}(L_{l})$, the previous statements still hold since $D_{l}$ and $B_{l}$ are closed sets. We now show that $B_{l}$ and $D_{l}$ satisfy Equation (16).
Proof of Equation (16). By applying Lemma 6 and its proof for arbitrary dimension $j$ to all $x\in{\rm conv}(L_{l})\cap U[I]$, there exist deathpoints $\overline{d}^{j}$ and $\underline{d}_{j}$ that satisfy $(\overline{d}^{j})_{j}=\sup_{d\in D_{l}}d_{j}$ and $(\underline{d}_{j})_{j}=\inf_{d\in D_{l}}d_{j}$ and $(\underline{d}_{j})_{j}\leq x_{j}\leq(\overline{d}^{j})_{j}$ for all $x\in{\rm conv}(L_{l})\cap U[I]$. Moreover, these points are located on lines in $L_{l}$ that are $\delta$-consecutive by definition. Thus, applying Lemma 2 $(ii)$ repeatedly on these pairs of lines for all dimensions, we end up with $D_{l}$ having a diagonal smaller than $\delta$. The same goes for birthpoints. ∎
5.1.2 Proof of Proposition 5
Proof.
Let $\mathbb{M}=\bigoplus_{i\in\mathcal{I}}{k^{I_{i}}}$ and $\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}=\bigoplus_{i\in\tilde{\mathcal{I}}}{k^{\tilde{I}_{i}}}$ be the interval decompositions of $\mathbb{M}$ and $\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}$, with induced matching functions $\sigma_{{\mathbb{M}}}$ and $\sigma$ respectively. In order to upper bound the bottleneck distance ${d_{\rm B}}({\mathbb{M}},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})$, one can upper bound the interleaving distance ${d_{\rm I}}({k^{I_{i}}},{k^{\tilde{I}_{\nu(i)}}})$ for any index $i\in\mathcal{I}$. Let $I$ and $\tilde{I}$ be two such intervals (we drop the index $i$ to alleviate notations). We need to show that the morphisms $f^{(\delta)}\colon{k^{I}}\rightarrow{k^{\tilde{I}}}(\boldsymbol{\delta})$ and $g^{(\delta)}\colon{k^{\tilde{I}}}\rightarrow{k^{I}}(\boldsymbol{\delta})$ exist and commute, i.e., that they induce a $\delta$-interleaving. Hence, we first show that:
$g^{(\delta)}_{x+\boldsymbol{\delta}}\circ f^{(\delta)}_{x}=\varphi^{x+2\boldsymbol{\delta}}_{x},$ (17)
for any $x\in K$, where $\varphi_{\cdot}^{\cdot}$ denote the transition maps of $k^{I}$.
If $x\in l$ for some line $l\in L$, Equation (17) is satisfied from ${I}\cap l={\tilde{I}}\cap l$, which itself comes from the fact that $\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}$ has the same $L$-fibered barcode than $\mathbb{M}$} (see Proposition 3 $(i)$). Hence, we assume in the following that $x\not\in\cup_{l\in L}l$. Furthermore, if $x\not\in{I}$ or $x+2\boldsymbol{\delta}\not\in{I}$, then Equation (17) is trivially satisfied. Hence, we also assume $x,x+2\boldsymbol{\delta}\in{I}$. This means that $b^{I}_{x}$ and $d^{I}_{x}$ are well-defined, and that $\varphi^{x+2\boldsymbol{\delta}}_{x}\cong\mathrm{id}_{k\to k}$. Thus we only have to show that ${(k^{\tilde{I}})}_{x+\boldsymbol{\delta}}\cong k$, i.e., $x+\boldsymbol{\delta}\in{\tilde{I}}$.
As $L=L_{\delta}(K^{\delta})$ is the $\delta$-grid of $K^{\delta}$ and $x\in K$, let $l\in L$ be a line such that $x\in{\rm conv}(L_{l})$ and let $l_{x}\subseteq{\rm conv}(L_{l})$ be the diagonal line passing through $x$. Now, as $R_{x,x+\boldsymbol{\delta}}\subseteq{I}$, Lemma 2 $(1)$ ensures that ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l}}\right)\neq\varnothing$ for any line $l\in L$ that is $\delta$-comparable to $l_{x}$; and the same holds for $\tilde{I}$ since ${\mathcal{F}}{\mathcal{B}}({k^{I}})_{L}={\mathcal{F}}{\mathcal{B}}({k^{\tilde{I}}})_{L}$. Using Proposition 7 on both $I$ and $\tilde{I}$, there exist two sets $B_{l}$ and $D_{l}$ such that $d_{x}^{I},d_{x}^{\tilde{I}}\in D_{l}$ and $b_{x}^{I},b_{x}^{\tilde{I}}\in B_{l}$, with the segments $l_{x}\cap B_{l}$ and $l_{x}\cap D_{l}$ having length at most $\delta$. Since one also has:
$b_{x}^{I}\leq x\leq x+2\boldsymbol{\delta}\leq d_{x}^{I},$
and $\left\lVert d_{x}^{I}-d_{x}^{\tilde{I}}\right\rVert_{\infty},\left\lVert b_{x}^{I}-b_{x}^{\tilde{I}}\right\rVert_{\infty}\leq\delta$, one finally has $b_{x}^{\tilde{I}}\leq x+\boldsymbol{\delta}\leq d_{x}^{\tilde{I}}$, which concludes that $x+\boldsymbol{\delta}\in{\tilde{I}}$. Since $I$ and $\tilde{I}$ are interchangeable in the arguments above, the result follows. ∎
5.2 Exact reconstruction
In this section, we identify the interval decomposable multi-parameter persistence modules that can be recovered exactly with our MMA algorithm. Given a precision parameter $\delta>0$, they correspond to modules that can decomposed into interval summands that form a subclass of the family of discretely presented interval modules, that we call the $\delta$-discretely presented interval modules.
Definition 24 ($\delta$-discretely presented interval module).
Let $K\subseteq\mathbb{R}^{n}$ be a rectangle in $\mathbb{R}^{n}$, and let $k^{I}$} be a discretely presented interval module. Given $\delta>0$, we say that $I$ is $\delta$-discretely presented in $K$ if:
  1. 1.
    (Large facets) for each point $x\in L[I]$ (resp. $U[I]$) there exists, for each facet $F$ containing $x$, an (n1)(n-1)-hypercube $Q^{x}_{F}$ of side length $2\delta$ such that $x\in Q^{x}_{F}$ and $Q^{x}_{F}\subseteq F$;
  2. 2.
    (Large holes) if there exists a diagonal line $l$ such that $l\cap I=\varnothing$, then there exists an $n$-hypercube $R$ of side length $\delta$ containing $\bf 0$} such that for any line $l^{\prime}$ in $l+R$, one has $l^{\prime}\cap I=\varnothing$;
  3. 3.
    (Locally small complexity) any $\infty$-ball of radius $\delta$, i.e., any set $B_{\delta}(x):=\{y\in\mathbb{R}^{n}:d_{\infty}(x,y)\leq\delta\}$ for some $x\in\mathbb{R}^{n}$, intersects at most one facet in $L[I]$ (resp. $U[I]$) of any fixed codirection;
  4. 4.
    (Compact description) each facet of $I$ has a non-empty intersection with $K$.
Assumptions 1 and 2 ensures that the parts of $I$ are large enough w.r.t. $\delta$, while Assumptions 3 and 4 ensure that surrounding sets of lines can detect at most one facet associated to a given codirection at a time, and that critical points of $I$ are all included in the rectangle $K$, respectively.
Remark 10.
One might wonder whether Assumption 2 and Assumption 3 are redundant with Assumption 1. In other words, one might wonder whether it is actually possible to define an interval module with large facets and small holes, or with large facets that can share the same codirection and lie close to each other at the same time. Even though this seems to be impossible when $n=2$ (indicating that Assumption 2 and Assumption 3 might indeed be redundant with Assumption 1 in that case), it can happen when $n\geq 3$, as Figure 11 shows.
Figure 11: Example of interval module in dimension $n=3$ with large facets, small holes and some facets with the same codirection close to each other. The support of the module can be constructed by taking the (closed) red and (open) green $L$-shaped sets on (Left), and glue them together as shown in (Middle). While arbitrarily large facets can be created using this construction, the resulting interval always contains a small hole and large facets of same codirection that are close to each other. Because of this, it is possible to find a (blue) diagonal line that goes through the support without intersecting it, while lines in its surrounding set will detect some facets. (Right) View of the interval from the top showing the hole and the spatially close facets (showed in bold font). This is an example where Assumptions 1 and 4 of Definition 24 are satisfied, while Assumptions 2 and 3 are not.
The main advantage of $\delta$-discretely presented modules is that they ensure that Algorithm 3 can identify every single facet with a corresponding label.
Lemma 8 (Labels are exact).
Let $\delta>0$ and $K$ be a rectangle in $\mathbb{R}^{n}$. Let $k^{I}$} be a $\delta$-discretely presented interval module in $K$, and let $L:=L_{\delta}(K^{2\delta})$ be the $\delta$-grid of lines of the offset $K^{2\delta}$. Then, there exists a bijection between the facets of $I$ and the labels identified by Algorithm 3.
Proof.
We first prove the result for birthpoints and facets of $L[I]$.
Let $F$ be a facet of $L[I]$. Let $l_{F}\in L$ be a diagonal line intersecting $F$, and $b_{F}\in\mathbb{R}^{n}$ be the associated birthpoint. By Definition 24, item (1), there exists an (n1)(n-1)-hypercube $Q_{F}^{b_{F}}\subseteq F$ of side length $2\delta$ such that $b_{F}\in Q_{F}^{b_{F}}$. This ensures that for any dimension $i$ that is not in the codirection: $i\in\llbracket 1,n\rrbracket\backslash{\rm codir}(F)$, one has either $b_{F}+\delta e_{i}\in Q_{F}^{b_{F}}$ or $b_{F}-\delta e_{i}\in Q_{F}^{b_{F}}$. Since $L$ is the $\delta$-grid of lines associated to $K^{2\delta}$, and since $Q_{F}^{b_{F}}$ is an (n1)(n-1)-hypercube, there exists a line $l_{0}\in L$ such that $l_{F}$ belongs to the surrounding set $L_{l_{0}}$, and such that the birthpoints corresponding to the lines in $L_{l_{0}}$ are all in $Q_{F}^{b_{F}}$. This means that ${\rm codir}(F)$ is detected as a label of $b_{F}$ by Algorithm 3.
Reciprocally, assume there exists a line $l_{0}\in L$ such that all birthpoints associated to the lines in the surrounding set $L_{l_{0}}$ share a coordinate along dimension $i\in\llbracket 1,n\rrbracket$, so that $i$ is a label detected by Algorithm 3. Then, the set of birthpoints $B_{L_{l_{0}}}$ has a minimal element, and thus its convex hull ${\rm conv}(B_{L_{l_{0}}})$ is in $L[I]$. Since ${\rm conv}(B_{L_{l_{0}}})$ is an (n1)(n-1)-hypercube of codirection $i$, it must be associated to a facet of $L[I]$ of codirection $i$ as well.
The proof extends straightforwardly for deathpoints.
Now that we have proved that all facets can be detected with $\delta$-grids of lines and $\delta$-discretely presented modules, we can state our following result, which claims that it is possible to exactly recover the underlying module under the same assumptions.
Lemma 9 (Exact recovery of intervals).
Let $\delta>0$ and $K=R_{\alpha,\beta}$ be a rectangle in $\mathbb{R}^{n}$, where $\alpha\leq_{n}\beta$. Let $k^{I}$} be a $\delta$-discretely presented interval module in $K$, and let $L:=L_{\delta}(K^{2\delta})$ be the $\delta$-grid of lines of the offset $K^{2\delta}$. Let $B=L\cap I$, and $C_{B}^{L}(B)$ and $C_{D}^{L}(B)$ be the $L$-birth and death corners of II computed by Algorithm 4, and let $\tilde{I}=\bigcup_{c\in C_{B}^{L}(B)}\bigcup_{c^{\prime}\in C_{D}^{L}(B)}R_{c,c^{\prime}}$ be the approximation computed by Algorithm 2. Then, one has:
${d_{\rm I}}\left({k^{I}},{k^{\tilde{I}}}\right)={d_{\rm B}}\left({k^{I}},{k^{\tilde{I}}}\right)=0.$ (18)
Proof.
As interval modules are characterized by their support, it is enough to show that $\overline{I}=\overline{\tilde{I}}$. In the following, we thus assume that $I$ is closed in $\overline{\mathbb{R}}^{n}$. We will also use an additional definition. Let $b$ be an infinite corner computed by Algorithm 4. We say that $b^{\prime}$ is a pseudo birth corner for $b$ if:
  1. 1.
    $b^{\prime}_{i}=b_{i}$ for all $i\in\llbracket 1,n\rrbracket\backslash\mathcal{J}$, and for each dimension $j\in\mathcal{J}$, there exists a hyperplane of codirection $j$ intersecting $K$ such that $\bigcap_{j}H_{j}\ni b^{\prime}$. The set $\mathcal{J}$ is called the codirection of $b^{\prime}$ and denoted with ${\rm codir}(b^{\prime})$, and the set $\llbracket 1,n\rrbracket\backslash\mathcal{J}$ is called the direction of $b^{\prime}$ and is denoted with ${\rm dir}(b^{\prime})$.
  2. 2.
    there exists a line $l_{0}\in L$ such that:
    1. (a)
      $b^{\prime}\in{\rm conv}(L_{l_{0}})\cap K^{2\delta}\backslash K$,
    2. (b)
      for each line $l\in L_{l_{0}}$, the endpoint $b_{l}^{I}$ is non trivial,
    3. (c)
      for each dimension $j\in\mathcal{J}$, there exists $l_{j}\in L_{l_{0}}$ such that $b_{l_{j}}^{I}\in H_{j}$.
Note that codirections can be extended to any finite corner (i.e., that is potentially not the pseudo corner of an infinite corner) straightforwardly, and that pseudo death corners can be defined by symmetry. We now prove Proposition 12.
We first show the inclusion ${\tilde{I}}\subseteq{I}$. More specifically, we have to prove that the corners computed by Algorithm 4 all belong to I{I}. A key argument that we will use several times comes from the following lemma, which allows for a local control of the boundary of I{I} using the hyperplanes associated to specific corners.
Lemma 10.
Let $b$ be a birthpoint (resp. deathpoint) of $I$ in $K^{\delta}$, and $l_{0}\in L$ be the line such that $b\in{\rm conv}(L_{l_{0}})$ (this line exists since $L$ fills $K^{\delta}$). Then, one has the following:
  1. 1.
    for any facet $F$ of $L[I]$ (resp. $U[I]$) containing $b$, there exists a line $l_{F}\in L_{l_{0}}$ such that $b_{l_{F}}^{I}\in F$ (resp. $d_{l_{F}}^{I}\in F$).
  2. 2.
    for any dimension $i$, there exists at most one facet of codirection $i$ intersecting the set of birthpoints (resp. deathpoints) $\left\{b_{l}^{I}:l\in L_{l_{0}}\right\}$ (resp. $\left\{d_{l}^{I}:l\in L_{l_{0}}\right\}$.
  3. 3.
    let $b^{\prime}_{L_{l_{0}}}$ (resp. $d^{\prime}_{L_{l_{0}}}$) be the finite corner generated by $L_{l_{0}}$. Then, one has:
    ${\rm conv}(L_{l_{0}})\cap L[I]\cap K^{2\delta}\subseteq\bigcup_{i\in{\rm codir}(b^{\prime})}\left\{x\in\mathbb{R}^{n}:x_{i}=b^{\prime}_{i}\right\}$
    (resp. ${\rm conv}(L_{l_{0}})\cap U[I]\cap K^{2\delta}\subseteq\bigcup_{i\in{\rm codir}(d^{\prime})}\left\{x\in\mathbb{R}^{n}:x_{i}=d^{\prime}_{i}\right\}\text{)}.$
Proof.
We only show the result for birthpoints since the arguments for deathpoints are the same. Let $b\in L[I]$ be a birthpoint in $K^{\delta}$.
Proof of (1). Let $F$ be a facet containing $b$. According to Definition 24, item (1), there exists an (n1)(n-1)-hypercube $Q^{b}_{F}$ of side length $2\delta$ such that $Q^{b}_{F}\subseteq F$ and $b\in Q^{b}_{F}$. Since $L$ is a grid, there exists a line $l\in L$ with $d_{\infty}(b,l)<\delta$ intersecting $Q^{b}_{F}$. Now, since $b\in{\rm conv}(L_{l_{0}})$, one has $d_{\infty}(l\cap H_{F},L_{l_{0}}\cap H_{F})<\delta$, where $H_{F}$ is the hyperplane containing $F$; thus, $l\in L_{l_{0}}$ (the argument is the same than in the proof of Proposition 7, first paragraph).
Proof of (2). By Proposition 7, item (2), the birthpoints associated to lines of $L_{l_{0}}$ are all contained in a ball of radius $\delta$. Thus, the unicity of the facets with given codirection comes straightforwardly from Definition 24, item (3).
Proof of (3). Note that the birthpoint $b$ is obviously included in the facets of $L[I]$ that contain it, which is a subset of the facets associated to the birthpoints of the lines in $L_{l_{0}}$. Now, as Lemma 8 ensures that the birthpoints associated to lines in $L_{l_{0}}$ are correctly labelled, the corner generated by $L_{l_{0}}$ must be on the intersection of the facets containing $b$. This ensures that:
$b\in\bigcup_{i\in\mathrm{codir}(b^{\prime})}\left\{x\in\mathbb{R}^{n}:x_{i}=b^{\prime}_{i}\right\}.$
Since these arguments do not depend on $b\in{\rm conv}(L_{l_{0}})$, the result follows. ∎
Now that we have Lemma 10, we can prove that finite and infinite corners belong to ${\rm supp}(I)$. We will prove the results for birth corners, but the arguments for death corners are symmetric.
Finite corners. Let $b$ be a finite birth corner, associated to a set of consecutive lines $L_{l_{0}}$ for some line $l_{0}\in L$. By assumption, each birthpoint $b_{l}^{I}$, for $l\in L_{l_{0}}$, is nontrivial; and thus any birthpoint in ${\rm conv}(L_{l_{0}})$ is nontrivial as well, using Definition 24, item (2). Let $l\in{\rm conv}(L_{l_{0}})$ be the diagonal line passing through $b$.
Using Lemma 10, one has:
$b_{l}^{I}\in{\rm conv}(L_{l_{0}})\cap L[I]\cap K^{2\delta}\subseteq\left(\bigcup_{i\in{\rm codir}(b)}\left\{x\in\mathbb{R}^{n}:x_{i}=b_{i}\right\}\right)\cap l=\left\{b\right\}.$
Thus $b=b_{l}^{I}$ and $b\in{I}$.
Infinite corners. Let $b$ be an infinite birth corner, and let $b^{\prime}$ be the minimal (w.r.t. $\leq_{n}$) pseudo birth corner for $b$, which is well defined by construction of $b$ (see Algorithm 4). Let $L_{l_{0}}$ be the associated set of consecutive lines $L_{l_{0}}$, for some line $l_{0}\in L$. We will show that, if $j$ is a free coordinate of $b^{\prime}$, i.e., if $j\in{\rm dir}(b^{\prime})$, then $b^{\prime}_{j}<\alpha_{j}$ (recall that $K$ is the rectangle $R_{\alpha,\beta}$). The reason we want to prove such inequalities is that they directly lead to the result. Indeed, if $b^{\prime}_{j}<\alpha_{j}$ for any $j\in{\rm dir}(b^{\prime})$, then $b^{\prime}-t\sum_{j\in{\rm dir}(b^{\prime})}e_{j}$ belongs to $L[I]$ for any $t>0$, since otherwise the line $\{b^{\prime}-t\sum_{j\in{\rm dir}(b^{\prime})}e_{j}:t>0\}$ would have to intersect a facet $F\subseteq L[I]$ of codirection $j$ for some $j\in{\rm dir}(b^{\prime})$, which would not intersect $K$, contradicting Definition 24, item (4).
Let $j\in{\rm dir}(b^{\prime})$ be a free coordinate. By contradiction, assume that $b^{\prime}_{j}\geq\alpha_{j}$, and let $b^{j}$ denote the pseudo birth corner generated by $L_{l_{0}-\delta e_{j}}$. In particular, this means that, for any $l\in L_{l_{0}}$, $l-\delta e_{j}\in L$ and $L_{l-\delta e_{j}}\subseteq L$ since $L$ fills $K^{\delta}$. Now, if for every line $l\in L_{l_{0}}$ such that $l=l_{0}+\overrightarrow{v}$ with $\overrightarrow{v}_{j}=0$, one has that $b_{l}^{I}$ and $b_{l-\delta e_{j}}^{I}$ are on the same facets, then one has $b_{l-\delta e_{j}}^{I}=b_{l}^{I}-\delta e_{j}$, and the pseudo corner $b^{j}$ is equal to $b^{\prime}-\delta e_{j}$ by construction, as per Algorithm 4. Moreover, one has $b^{j}=b^{\prime}-\delta e_{j}\leq_{n}b^{\prime}$, contradicting the fact that $b^{\prime}$ is minimal. Hence, there is at least one line $\in L_{l_{0}}, l=l$$l_{0}+\overrightarrow{v}$ with $\overrightarrow{v}_{j}=0$, such that $b_{l}^{I}$ and $b_{l-\delta e_{j}}^{I}$ are not on the same facets, in other words, there exists a facet $F_{j}$ of $L[I]$ of codirection $j$ that intersects the (half-open) segment $b_{l}^{I}-\delta e_{j},b_{l}^{I})$. In order to locate that facet more precisely, we will prove the following lemma.
Lemma 11.
For any $i\in\llbracket 1,n\rrbracket$ and $s,t\in\mathbb{R}$ such that $s<t$, one has $(b^{I}_{l-te_{i}})_{i}\leq(b^{I}_{l-se_{i}})_{i}$.
Proof.
Without loss of generality, assume $s=0$. Since $b^{I}_{l}-te_{i}\in l-te_{i}$, it follows that $b^{I}_{l}-te_{i}$ and $b^{I}_{l-te_{i}}$ are comparable. Moreover, one must have $b^{I}_{l}-te_{i}\leq_{n}b^{I}_{l-te_{i}}$, otherwise one would have $b^{I}_{l}>_{n}b^{I}_{l}-te_{i}>_{n}b^{I}_{l-te_{i}}$, contradicting Lemma 1. If the points are equal, i.e., $b^{I}_{l}-te_{i}=b^{I}_{l-te_{i}}$, then one has $(b^{I}_{l})_{i}\geq(b^{I}_{l-te_{i}})_{i}$. Otherwise, if $b^{I}_{l}-te_{i}<_{n}b^{I}_{l-te_{i}}$, then:
$\forall k\neq i,\,(b^{I}_{l-te_{i}})_{k}>(b^{I}_{l})_{k}.$
Moreover, since $b^{I}_{l}$ and $b^{I}_{l-te_{i}}$ cannot be comparable as per Lemma 1 one must have $(b^{I}_{l-te_{i}})_{i}\leq(b^{I}_{l})_{i}$.
Let $H_{j}=\left\{x\in\mathbb{R}^{n}:x_{j}=c_{j}\right\}$ be the hyperplane associated to $F_{j}$. Then, by Lemma 11, one has:
$(b_{l-\delta e_{j}}^{I})_{j}\leq c_{j}<(b_{l}^{I})_{j}.$
Since the lines $l$ and $l-\delta e_{j}$ both belong to the surrounding set $L_{l_{0}-\delta e_{j}}$, it follows from Lemmas 8, and 10, item (3), that ${\rm codir}(b^{j})\supseteq{\rm codir}(b^{\prime})\cup\left\{j\right\}$. Moreover, since the facets of $L[I]$ associated to ${\rm codir}(b^{j})$ are unique in a $\delta$-ball around $b^{j}$, as per Definition 24, item (3), they all have a unique associated value $c_{i}$ (corresponding to their associated hyperplanes).
Finally, we will show that $b^{j}\leq_{n}b^{\prime}$. Let $i\in\llbracket 1,n\rrbracket$ be an arbitrary dimension.
  • If $i\in{\rm codir}(b^{\prime})$, then $b^{j}_{i}=b^{\prime}_{i}$.
  • If $i\in{\rm codir}(b^{j})\backslash{\rm codir}(b^{\prime})$, then $b^{j}_{i}\in\left\{c_{i},\min_{l\in L_{l_{0}-\delta e_{j}}}(b_{l}^{I})_{i}\right\}\leq\min_{l\in L_{l_{0}}}(b_{l}^{I})_{i}=b^{\prime}_{i}$, with a strict inequality for $i=j$.
  • If $i\in{\rm dir}(b^{j})\subseteq{\rm dir}(b^{\prime})$, then $b^{j}_{i}=\min_{l\in L_{l_{0}-\delta e_{j}}}(b_{l}^{I})_{i}\leq\min_{l\in L_{l_{0}}}(b_{l}^{I})_{i}=b^{\prime}_{i}$.
Hence, one always has $b^{j}_{i}\leq b^{\prime}_{i}$, and thus $b^{j}<_{n}b^{\prime}$, which contradicts the fact that $b^{\prime}$ is minimal. Thus, one must have $b^{\prime}_{j}<\alpha_{j}$.
We now show that ${I}\subseteq{\tilde{I}}$. Let $x\in{I}$. We will show that there exists a birth corner $c$ such that $c\leq_{n}x$. Let $\mathcal{H}$ be the family of hyperplanes associated to the facets of $L[I]$. The corner $c$ will be defined as the limit of a sequence of points $\{x^{(k)}\}_{k\in\mathbb{N}^{*}}$in $\overline{\mathbb{R}}^{n}$, defined by induction with:
  1. 1.
    $x^{(1)}=\inf\left\{x-t\cdot\mathbf{1}:t\geq 0\right\}\cap{\rm supp}(I)$. Then, one has the two following possibilities:
[…]
    • or there exists a maximal subset of hyperplanes $\mathcal{H}^{1}\subset\mathcal{H}$, $\mathcal{H}^{1}\neq\varnothing$, such that $x^{(1)}\in\cap_{H\in\mathcal{H}^{1}}H=:H_{1}$. Let $\mathcal{J}^{1}\subseteq\llbracket 1,n\rrbracket$ be the set of free coordinates in $H_{1}$, i.e., those dimensions such that $j\in\mathcal{J}^{1}\Longleftrightarrow x^{(1)}-e_{j}\in H_{1}$.
  2. 2.
    $x^{(2)}=\inf\left\{x^{(1)}-t\cdot\sum_{j\in\mathcal{J}^{1}}e_{j}:t\geq 0\right\}\cap{\rm supp}(I)$. Then, one has the two following possibilities:
    • either $x^{(2)}$ is at infinity in $H_{1}$, i.e., $x^{(2)}_{j}$$\infty$ if $j\in\mathcal{J}^{1}$ and $x^{(2)}_{j}=x^{(1)}_{j}$ otherwise, and we let $c:=x^{(2)}$.
    • or there exists a maximal subset of hyperplanes $\mathcal{H}^{2}\supsetneq\mathcal{H}^{1}$ such that $x^{(2)}\in\cap_{H\in\mathcal{H}^{2}}H=:H_{2}$. Let $\mathcal{J}^{2}\subseteq\llbracket 1,n\rrbracket$ be the set of free coordinates in $H_{2}$, i.e., those dimensions such that $j\in\mathcal{J}^{2}\Longleftrightarrow x^{(2)}-e_{j}\in H_{2}$.
  3. 3.
    For $k\geq 3$, $x^{(k+1)}=\inf\left\{x^{(k)}-t\cdot\sum_{j\in\mathcal{J}^{k}}e_{j}:t\geq 0\right\}\cap{\rm supp}(I)$. Then, one has the two following possibilities:
    • either $x^{(k+1)}$ is at infinity in $H_{k}$, i.e., $x^{(k+1)}_{j}$$\infty$ if $j\in\mathcal{J}^{k}$ and $x^{(k+1)}_{j}=x^{(k)}_{j}$ otherwise, and we let $c:=x^{(k+1)}$.
    • or there exists a maximal subset of hyperplanes $\mathcal{H}^{k+1}\supsetneq\mathcal{H}^{k}$ such that $x^{(k+1)}\in\cap_{H\in\mathcal{H}^{k+1}}H=:H_{k+1}$. Let $\mathcal{J}^{k+1}\subseteq\llbracket 1,n\rrbracket$ be the set of free coordinates in $H_{k+1}$, i.e., those dimensions such that $j\in\mathcal{J}^{k+1}\Longleftrightarrow x^{(k+1)}-e_{j}\in H_{k+1}$.
[…]
$c\leq_{n}c^{-}\leq_{n}\cdots\leq_{n}x^{(2)}\leq_{n}x^{(1)}\leq_{n}x.$
We now show that $c$ is indeed a birth corner. If $c$ is finite, then it must belong to the intersection of $n$ hyperplanes, and it is thus a finite birth corner. Hence, we assume now that $c$ is not finite. We will construct a minimal pseudo birth corner from $c^{-}$, and show that $c$ is its associated infinite birth corner. We will consider two different cases, depending on whether $c^{-}$ is close to $K=R_{\alpha,\beta}$ or not. If $c^{-}\in K^{\delta}$, the filling property of $L$ and the size of the facets of $L[I]$ ensure that $c^{-}$ is itself a minimal pseudo birth corner, associated to $c$, which is thus an infinite birth corner. If $c^{-}\notin K^{\delta}$, then let $\overrightarrow{v}\in\mathbb{R}^{n}$ be a vector that pushes back $c^{-}$ into $K^{\delta}$, i.e., such that, for any dimension $i\in\mathcal{J}^{-}$, one has:
$\alpha_{i}-\delta\leq(c^{-}+\overrightarrow{v})_{i}<\alpha_{i},$
and $\overrightarrow{v}_{i}=0$ if $i\notin\mathcal{J}^{-}$. Let $S$ be the segment $c^{-},c^{-}+\overrightarrow{v}]$. We have the two following cases:
  1. 1.
    Assume $S\subseteq L[I]$. Then $c^{-}+\overrightarrow{v}\in{\rm supp}(I)\cap K^{\delta}$, and there exists a line $l\in L$ such that $c^{-}+\overrightarrow{v}\in{\rm conv}(L_{l})$. Let $c^{l}$ be the pseudo birth corner associated to $L_{l}$. Since one has $c^{l}_{j}<\alpha_{j}$ for any dimension $j\in\mathcal{J}^{-}$, it follows that $\mathcal{J}^{-}\subseteq{\rm dir}(c^{l})$. Furthermore, since $c^{-}+\overrightarrow{v}$ belongs to the same facets than $c$ and $c^{-}$, and since $c^{-}+\overrightarrow{v}\in{\rm conv}(L_{l})$ one has ${\rm codir}(c^{l})\supseteq{\rm codir}(c)$ and ${\rm dir}(c)=\mathcal{J}^{-}$. Thus, $c$ is an infinite birth corner associated to the minimal pseudo birth corner $c^{l}$.
  2. 2.
    Assume $S\not\subseteq L[I]$. In that case, there must be a facet of codirection $j$, for some $j\in\mathcal{J}^{-}$, that intersects $S$. Since one has $c^{-}_{j}\leq(c^{-}+\overrightarrow{v})_{j}<\alpha_{j}$ for any $j\in\mathcal{J}^{-}$, this means that the facet would not intersect $K$, which yields to a contradiction as per Definition 24, item (4).
This concludes that ${I}\subseteq{\tilde{I}}$, and the equality between these supports holds. ∎
Lemma 9 extends to the following proposition, whose proof is immediate from the definition of induced matchings.
Proposition 12 (Exact recovery).
Let $\mathbb{M}$ be a f.p. interval decomposable $n$-parameter persistence module. Let $K$ be a rectangle in $\mathbb{R}^{n}$ that compactly characterizes $\mathbb{M}$, and $L:=L_{\delta}(K^{2\delta})$ be the $\delta$-grid of lines of the offset $K^{2\delta}$. Assume that all interval summands of $\mathbb{M}$ are $\delta$-discretely presented, and let $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}:=\mathtt{MMA}{}({\mathbb{M}},L,\sigma)$}, where $\sigma$ is a matching function that commutes with the induced matching function $\sigma_{\mathbb{M}}$. Then, one has:
${d_{\rm I}}({\mathbb{M}},\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta})={d_{\rm B}}({\mathbb{M}},\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta})=0.$
Note that f.p. interval decomposable modules are always made of $\delta$-discretely presented interval summands, for small enough yet positive $\delta$ (one can take for instance the smallest distance $\delta_{\rm exact}>0$ between two distinct graded Betti numbers).
One might wonder whether the usual distances between barcodes, such as the bottleneck or Wasserstein distances, could be used to define matching functions that commute with induced matching functions. Indeed, a major advantage of, e.g., Wasserstein distances, is that their associated matching functions are usually unique. However, when the space $\delta$ between two lines is too large, the matching functions induced by Wasserstein distances can still fail to be induced, as shown in Figure 12. In the next section, we discuss how to design such matching functions from compatible matching functions.
Figure 12: Example of interval decomposable module with two interval summands (green and purple), and its barcodes along two lines (here the two couples of red-blue bars). Any matching function induced by, e.g., Wasserstein distances between the barcodes, will match the first red bar with the second red bar and the first blue bar with the second blue bar; however, this matching is not induced.
6 Finding compatible and induced matching functions
In this section, we discuss how to design matching functions that are compatible or that commute with induced matching functions, so as to satisfy the assumptions of Proposition 3, Proposition 5 and Proposition 12. In particular, in Section 6.1 we restrict to interval decomposable modules, and we show that compatible and usual matching functions always commute with induced matching functions for small enough $\delta$ and generic assumptions. Then, in Section 6.2, we show that the vineyards algorithm induces compatible matching functions when $n=2$.
6.1 Induced matching functions for interval decomposable modules
In this section, we show, given some interval decomposable module $\mathbb{M}$, that compatible matching functions commute with the induced matching function $\sigma_{{\mathbb{M}}}$ under some specific conditions. In the rest of this section, given a matching function $\sigma$ that commutes with $\sigma_{\mathbb{M}}$, we will also call $\sigma$ induced for the sake of simplicity. Before stating the main result, we first show a technical lemma that relates the locations of endpoints of compatible bars with each other.
Lemma 13.
Let $l_{1}$ and $l_{2}$ be two $\delta$-consecutive lines, and let $b_{1},d_{1}):={\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)$ be the bar of a f.p. interval module along $l_{1}$. Let $b_{2},d_{2})$ be a bar along $l_{2}$ that is compatible with $b_{1},d_{1})$. Then, $d_{2}$ (resp. $b_{2}$) is included in a segment of size $\delta$ in $l_{2}$.
Proof.
Applying Lemma 2, one has:
$d_{2}\in C:=\big[B_{\delta}(d_{1})\cap l_{2}\big]\backslash\big[\left\{z\in\mathbb{R}^{n}:z>d_{1}\right\}\cup\left\{z\in\mathbb{R}^{n}:z<d_{1}\right\}\big].$
Since $C$ is a nonempty, totally ordered set, we can define $y:=\min C$. By construction, there exists a dimension $i$ such that $y_{i}\geq(d_{1})_{i}$, and thus $C$ must be included in the segment $y,y+\delta\cdot\mathbf{1}]$ along $l_{2}$.
The proof applies straightforwardly to $b_{2}$ by symmetry. ∎
Since bars that are matched under an induced matching function are always compatible, one way to construct an induced matching function between two barcodes is therefore to isolate, among all possible matching functions, the ones such that matched bars are compatible. If this family contains a single element, it must be the induced matching we are looking for. This typically happens for interval decomposable multi-parameter persistence modules whose summands are sufficiently separared, as we show in the proposition below.
Proposition 14.
Let ${\mathbb{M}}=\bigoplus_{I\in\mathcal{I}}{k^{I}}$ be a f.p. interval decomposable $n$-parameter persistence module. Let $\delta>0$, and ${k^{I}},{k^{I^{\prime}}}$ be two interval summands in the decomposition of $\mathbb{M}$}. Assume that the two following properties are satisfied:
  1. 1.
    Let $l\subset\mathbb{R}^{n}$ be a diagonal line such that ${I}\cap l\neq\varnothing$ and ${I^{\prime}}\cap l\neq\varnothing$.
    Then, one has either $\left\lVert b_{l}^{I}-b_{l}^{I^{\prime}}\right\rVert_{\infty}>\delta$ or $\left\lVert d_{l}^{I}-d_{l}^{I^{\prime}}\right\rVert_{\infty}>\delta$. In other words, the endpoints of the bar in ${\mathcal{B}}\left({{\left.\kern-1.2ptk^{I}\vphantom{\big|}\right|_{l}}}\right)$ and of the bar in ${\mathcal{B}}\left({{\left.\kern-1.2ptk^{I^{\prime}}\vphantom{\big|}\right|_{l}}}\right)$ are at distance at least $\delta$.
  2. 2.
    The bars of length at most $2\delta$ in $I$ and $I^{\prime}$ are at distance at least $\delta$, i.e., if we let:
    $S^{I}:=\left\{l:l\cap{I}\neq\varnothing,\left\lVert b_{l}^{I}-d_{l}^{I}\right\rVert_{\infty}\leq 2\delta\right\},$
    (and similarly for $I^{\prime}$), one has
    $d_{\infty}\left(S^{I},S^{I^{\prime}}\right)>\delta/2.$
    In other words, a small bar in $I$ cannot be too close to a small bar in $I^{\prime}$.
Then, the matching function $\sigma_{\rm comp}$, induced by matching bars that are compatible together, is well-defined and induced from $\mathbb{M}$.
Note that, upon using chunk reduction (36) and infinitesimal perturbations, or whenever the graded Betti numbers of $\mathbb{M}$ are independent and identically distributed from a non-singular distribution, it is always possible to ensure that Assumptions (1) and (2) are satisfied for a given f.p. interval decomposable module $\mathbb{M}$ and small enough $\delta$ (see also the paragraph after Proposition 12). See Figure 13 for an illustration of Assumptions (1) and (2).
Figure 13: (Left) Example of module whose interval summands do not satisfy Assumption (2). (Right) Example of module whose interval summands do satisfy Assumptions (1) and (2). Bars corresponding to consecutive lines can only be matched if they are compatible, which, in this figure, means that they have the same color, i.e., that they are associated to the same interval summand.
Proof.
Let $k^{I}$ and $k^{I^{\prime}}$ be two interval summands in the decomposition of $\mathbb{M}$}. Let $l_{1}$ and $l_{2}$ be two $\delta$-consecutive lines of $L$, and let $b:={\mathcal{B}}\left({\left.\kern-1.2ptk^{I}\vphantom{\big|}\right|_{l_{1}}}\right)$ be the bar corresponding to $I$ along $l_{1}$. We will show that $\sigma_{\rm comp}$ must match $b$ to either $b^{\prime}:={\mathcal{B}}\left({\left.\kern-1.2ptk^{I}\vphantom{\big|}\right|_{l_{2}}}\right)$ if ${I}\cap l_{2}\neq\varnothing$, or the empty set if ${I}\cap l_{2}=\varnothing$.
  • $\bullet$
    If ${\rm supp}(I)\cap l_{2}=\varnothing$, then by Lemma 2, the length of $b$ is at most $\delta$, i.e., $\left\lVert b^{I}_{l_{1}}-d^{I}_{l_{1}}\right\rVert_{\infty}\leq\delta$.
    It is thus compatible with the empty set. Now, since $d_{\infty}(l_{1},l_{2})=\delta/2$ and since $l_{1}\in S^{I}$, Assumption (2) ensures that the bar $b^{\prime\prime}:={\mathcal{B}}\left({\left.\kern-1.2ptk^{I^{\prime}}\vphantom{\big|}\right|_{l_{2}}}\right)$ (if it exists) must be of length at least $2\delta$. In particular, it is not compatible with $b$, hence $\sigma_{\rm comp}$ cannot match $b$ to $b^{\prime\prime}$, and must match $b$ to the empty set.
  • $\bullet$
    If ${\rm supp}(I)\cap l_{2}\neq\varnothing$, then the bar $b^{\prime}=[b^{I}_{l_{2}},d^{I}_{l_{2}})$ in ${\mathcal{B}}\left({\left.\kern-1.2ptk^{I}\vphantom{\big|}\right|_{l_{2}}}\right)$
    is compatible with $b$, as per Lemma 1. According to Lemma 13, it follows that the birthpoint and deathpoint of any bar along $l_{2}$ that is compatible to $l_{1}$ must belong to segments $s_{b},s_{d}$ of length $\delta$ that contain $b^{I}_{l_{2}}$ and $d^{I}_{l_{2}}$ respectively. Let $b^{\prime\prime}:=[b^{I^{\prime}}_{l_{2}},d^{I^{\prime}}_{l_{2}})$ be the bar in ${\mathcal{B}}\left({\left.\kern-1.2ptk^{I^{\prime}}\vphantom{\big|}\right|_{l_{2}}}\right)$ (if it exists).
    According to Assumption (1), we either have $\left\lVert b_{l_{2}}^{I}-b_{l_{2}}^{I^{\prime}}\right\rVert_{\infty}>\delta$ or $\left\lVert d_{l_{2}}^{I^{\prime}}-d_{l_{2}}^{I^{\prime}}\right\rVert_{\infty}>\delta$. In particular this means that either $b^{I^{\prime}}_{l_{2}}\not\in s_{b}$ or $d^{I^{\prime}}_{l_{2}}\not\in s_{d}$. Hence $b^{\prime\prime}$ is not compatible with $b$, and $\sigma_{\rm comp}$ must match $b$ to $b^{\prime}$.
In both cases, $\sigma_{\rm comp}$ is well-defined and induced from $\mathbb{M}$. ∎
One can check that the proof of Proposition 14 extends easily to the matching functions associated to the Wasserstein distances and the vineyards algorithm. Indeed, their associated matching functions are unique when $\delta$ becomes small enough, and thus must correspond to the only compatible matching $\sigma_{\rm comp}$ identified in Proposition 14.
6.2 The vineyards algorithm for general $2$-parameter modules
In this section, we show that the matching function associated to the vineyards algorithm for simplicial complexes is compatible, for small enough $\delta$ and $n=2$. This section is quite technical and can be skipped by readers who are most interested in the general exposition. Since vineyards are heavily based on simplicial homology, we first recall the basics of persistent homology from simplicial complexes in Section 6.2.1. Then, we provide an analysis of the vineyards algorithm in Section 6.2.2.
6.2.1 Persistent homology of simplicial complexes
We assume in the following that the reader is familiar with simplicial complexes, boundary operators and homology groups, and we refer the interested reader to (53, Chapter 1) for a thorough treatment of these notions. The first important definition is the one of filtered simplicial chain complexes.
Definition 25.
Let $S$ be a simplicial complex, and $f\colon S\to\mathbb{R}$ be a filtration function, i.e., $f$ satisfies $f(\sigma)\leq f(\tau)$ when $\sigma\subseteq\tau$. Then, the filtered simplicial chain complex $(S,f)$ is defined as $(S,f)=((C_{t})_{t\in\mathbb{R}},\iota)$, where:
  1. 1.
    $C_{t}=\left<\sigma_{0},\dots,\sigma_{i}\right$> is the vector space over a field $k$ whose basis elements are the simplices that have filtration values smaller than $t$, i.e., $\left\{\sigma_{0},\dots,\sigma_{i}\right\}=\left\{\sigma\in S:f(\sigma)\leq t\right\}$, and
  2. 2.
    for any $s\leq t$, the map $\iota=\iota_{s}^{t}\colon C_{s}\hookrightarrow C_{t}$ is the canonical injection.
Note that $f$ can be used to define an order on the simplices of $S=\left\{\sigma_{i}\right\}_{i=0}^{N}$, by using the ordering induced by the filtration values. In other words, we assume in the following that $f(\sigma_{0})\leq f(\sigma_{1})\leq\dots\leq f(\sigma_{N})$. We also slightly abuse notations and define $C_{i}:=\left<\sigma_{0},\dots,\sigma_{i}\right$> for any $i\in\llbracket 0,N\rrbracket$, and
[…]
Then, applying the homology functor $H_{*}$ on this filtered simplicial chain complex yields the following (single-parameter) persistence module:
$H_{*}(S,f)=0\rightarrow H_{*}(C_{0})\rightarrow H_{*}(C_{1})\rightarrow\dots\rightarrow H_{*}(C_{N}).$
An important theorem of single-parameter persistent homology states that, up to a change of basis, it is possible to pair some chains together in order to define the so-called one-dimensional persistence barcode associated to the filtered simplicial chain complex.
Theorem 15 (Persistence pairing, (30, Theorem 2.6)).
Given a filtered simplicial chain complex $(S,f)=C_{1}\hookrightarrow C_{2}\hookrightarrow\dots\hookrightarrow C_{N}$ and associated persistence module $H_{*}(S,f)$, there exists a partition $\llbracket 1,N\rrbracket=E\sqcup B\sqcup D$, a bijective map ${\rm Low}:D\rightarrow B$, and a new basis $\hat{\sigma}_{1},\dots,\hat{\sigma}_{N}$ of $C$, called reduced basis, such that:
  1. 1.
    $C_{i}=\langle\hat{\sigma}_{1},\dots,\hat{\sigma}_{i}\rangle$,
  2. 2.
    $\partial\hat{\sigma}_{e}=0$ for any $e\in E$,
  3. 3.
    for any $d\in D$, one has $\partial\hat{\sigma}_{{\rm Low}(d)}=0$, and $\partial\hat{\sigma}_{d}$ is equal to $\hat{\sigma}_{{\rm Low}(d)}$ up to simplification, i.e., there exists a set of indices ${\rm bd}(d)$ such that $(i)$ $j<{\rm Low}(d)\leq d$ for any $j\in{\rm bd}(d)$, and $(ii)$ $\partial\hat{\sigma}_{d}=\hat{\sigma}_{{\rm Low}(d)}+\sum_{j\in{\rm bd}(d)}\hat{\sigma}_{j}$.
In particular, the chains $\{\hat{\sigma}_{j}:j\in E\cap\llbracket 1,i\rrbracket\}\cup\{\hat{\sigma}_{j}:j\in B\cap\llbracket 1,i\rrbracket\text{ and }\exists d>i\text{ s.t. }{\rm Low}(d)=j\}$ form a basis of the simplicial homology groups $H_{*}(C_{i})$. Moreover, the chains $\{\hat{\sigma}_{j}:j\in B\sqcup E\}$ are called positive chains while the chains $\{\hat{\sigma}_{j}:j\in D\}$ are called negative chains.
The multiset of bars $\mathcal{B}(f):=\{[f(\hat{\sigma}_{b}),f(\hat{\sigma}_{d})]:b={\rm Low}(d)\}\cup\{[f(\hat{\sigma}_{e}),+\infty):e\in E\}$ is called the persistence barcode of the filtered simplicial chain complex $(S,f)$ and of the single-parameter persistence module $H_{*}(S,f)$.
Note that while the reduced basis $\{\hat{\sigma}_{1},\dots,\hat{\sigma}_{N}\}$ does not need to be unique, the pairing map $\rm Low$} is actually independent of that reduced basis, see (35, VII.1, Pairing Lemma).
6.2.2 Vineyards algorithm and matching
The vineyards algorithm (28) is a method that allows to find reduced chain bases for filtered simplicial complexes whose simplex orderings only differ by a single transposition of consecutive simplices, that we denote by (ii+1)(i\,i+1). This algorithm was later generalized to the setup of zigzag persistence modules in (52). This is the setup that we use in our context as follows. We start with some $n$-parameter persistence module $\mathbb{M}$, such that there exists a finite dimensional, generic, $n$-filtered one-critical simplicial chain complex $(S,F)$, satisfying $\mathbb{M}=H_{*}(F)$. [^5] We drop the dependence on $S$ for simplicity. We also fix an order $(\sigma_{1},\dots,\sigma_{N})$ on the simplices of $S$, and use the following notation (in this section only):
$(x_{1},\dots,x_{N}):=\left(F(\sigma_{1}),\dots,F(\sigma_{N})\right),\quad\textnormal{and }\quad(x_{1}^{l},\dots,x_{N}^{l}):=\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}(\sigma_{1}),\dots,{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}(\sigma_{N})\right),$ (20)
the filtration values (in $\mathbb{R}^{n}$) of the simplices of $F$ and $\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}$} respectively, for any given positive line $l$. Note that we always have, for any line $l$:
$\forall 1\leq j\leq N,\quad x_{j}\leq_{n}x_{j}^{l}\quad\textnormal{and}\quad\exists 1\leq k\leq n,\quad(x_{j})_{k}=(x_{j}^{l})_{k}.$ (21)
Finally, we consider the map $\mathrm{ord}\colon{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}\mapsto\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}\right)\in\mathfrak{S}_{N}$ that gives the partial order of the simplices of $F$ according to the filtration $\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}$}, completed to a total order by the initial order if necessary, i.e., $\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}\right)$ is the only permutation $\gamma\in\mathfrak{S}_{N}$ satisfying
[…]
Proposition 16.
Consider $l_{1},l_{2}$ two consecutive diagonal lines such that for any diagonal line $l_{1}\leq l\leq l_{2}$, we have either $\gamma=\gamma_{1}$ or $\gamma=\gamma_{2}$, where $\gamma:=\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l}}\right),\,\gamma_{1}:=\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}\right)$ and $\gamma_{2}:=\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}\right)$. Then, either $\gamma_{1}=\gamma_{2}$ or there exists an integer $1\leq i\leq n$ such that $\gamma_{1}=\gamma_{2}\circ(i\,i+1)$, and the $i$th vineyard update induces a compatible matching between ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{1}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{2}}}\right)$.
Remark 11 (More details about the vineyard matching).
More specifically, the matching function given by the vineyards algorithm corresponds to the following. If $(\tau_{1},\dots,\tau_{N})$ is a reduced basis on the first line, then this algorithm provides an updated basis $(\mathrm{vine}_{i}(\tau_{1}),\dots,\mathrm{vine}_{i}(\tau_{N}))$ that is reduced on the second line. The matching function is then given by the following relation: if $b$ (resp. $d$) is the birthpoint (resp. deathpoint) of a bar in the barcode induced by $\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{1}}$}, and generated by some $\tau_{j}$ (resp. $\tau_{k}$), then the endpoint $b$ is matched to the endpoint generated by $\mathrm{vine}_{i}(\tau_{j})$ (resp. $\mathrm{vine}_{i}(\tau_{k})$).
Note that a non-trivial bar $\left[{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{j}),{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{k})\right)$ may be matched to an empty bar if ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}((\tau_{j}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}(\tau_{k}))$. In particular, such continuous matching functions lead to indicator summands, which, once split up, lead to the desired candidate decomposition with interval summands.
Proof.
First note that if $\gamma_{1}=\gamma_{2}$, there is nothing to show, hence we assume $\gamma_{1}\neq\gamma_{2}$ in the following. Without loss of generality, we will first assume that the initial order $(\sigma_{1},\dots,\sigma_{N})$ on the first line is compatible with the partial order defined on the first line, i.e., $\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}\right)=\mathrm{id}$.
First, note that on the set of diagonal lines in $\mathbb{R}^{n}$, the map $l\mapsto x^{l}$ is continuous, and the same goes for the map $l\mapsto(x_{\pi_{1}}^{l},\dots,x_{\pi_{N}}^{l})$, where $\pi=\mathrm{ord}\left({\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}\right)$. In particular, if $\gamma_{1}\neq\gamma_{2}$ then there exists an integer $i$ such that $x_{\pi_{i}}^{l}=x_{\pi_{i+1}}^{l}$ for some line $l_{1}\leq l\leq l_{2}$. By the genericity assumption, this $i$ is unique which concludes that $\gamma_{1}=\gamma_{2}\circ(i\,i+1)$.
Now, as $\gamma_{1}\neq\gamma_{2}$, we have $x_{i}^{l_{1}}\leq x_{i+1}^{l_{1}}$ and $x_{i}^{l_{2}}>x_{i+1}^{l_{2}}$. Note that in particular that this implies that both $x_{i}$, $x_{i+1}$ and $x_{i}^{l_{1}}$, $x_{i+1}^{l_{2}}$ are strictly incomparable:
  • as $x_{i}^{l_{2}}>x_{i+1}^{l_{2}}$, there exists a dimension $j$ (obtained with Equation (21)) such that $\left(x_{i}\right)_{j}=\left(x_{i}^{l_{2}}\right)_{j}>\left(x_{i+1}^{l_{2}}\right)_{j}\geq\left(x_{i+1}\right)_{j}$}, and
  • If $x_{i}$ and $x_{i+1}$ were comparable, we would have (by previous point) $x_{i+1}\leq_{n}x_{i}$ with $x_{i}\neq x_{i+1}$, which would contradict the initial assumption, i.e., the fact that the initial order $\left(\sigma_{1},\dots,\sigma_{N}\right)$ is given by a completion of the original poset. Hence, there exists another dimension $j^{\prime}\neq j$ (obtained again with Equation (21)) such that $\left(x_{i}\right)_{j^{\prime}}\leq\left(x_{i}^{l_{1}}\right)_{j^{\prime}}<\left(x_{i+1}^{l_{1}}\right)_{j^{\prime}}=\left(x_{i+1}\right)_{j^{\prime}}$.
Combining everything, we have:
$\exists 1\leq j\neq j^{\prime}\leq n,\quad\left(x_{i}^{l_{1}}\right)_{j^{\prime}}<\left(x_{i+1}\right)_{j^{\prime}}\leq\left(x_{i+1}^{l_{2}}\right)_{j^{\prime}}\quad\textnormal{ and }\quad\left(x_{i+1}^{l_{2}}\right)_{j}<\left(x_{i}\right)_{j}\leq\left(x_{i}^{l_{1}}\right)_{j},$ (23)
which guarantees that $x_{i}^{l_{1}}$ and $x_{i+1}^{l_{2}}$ are strictly incomparable.
Consider for any index $1\leq j\leq N$ the complex $K_{j}:=\left<\sigma_{1},\dots\sigma_{j}\right$>, and the following diamond
$H_{*}\left(K_{i-1}\cup\left\{\sigma_{i+1}\right\}\right)$$\cdots$$H_{*}(K_{i-1})$$H_{*}\left(K_{i-1}\cup\left\{\sigma_{i}\right\}\right)$$H_{*}(K_{i+1})$$\cdots$$\scriptstyle{d}$$\scriptstyle{a}$$\scriptstyle{b}$$\scriptstyle{c}$ (24)
where the maps $a,b,c,d$ are induced by the inclusion. Note that as simplices are only added one by one, the maps $a,b,c,d$ are either surjerctive of nullity one or injective of corank one, depending on the positivity or negativity of the simplices $\sigma_{i}$ and $\sigma_{i+1}$. Furthermore, the Mayer-Vietoris theorem ensures that the following sequence is exact:
$H_{*}(K_{i-1})\xrightarrow{x\mapsto(x,x)}{H_{*}(K_{i-1}\cup\{\sigma_{i+1}\})\oplus K_{i-1}\cup\{\sigma_{i}\})}\xrightarrow{(x,y)\mapsto y-x}{H_{*}(K_{i+1})}.$
Note that in the continous case, Equation 24 corresponds to the following diagram:
$H_{*}\left(F_{x_{i+1}^{l_{2}}}\right)$$H_{*}\left(F_{x_{i}^{l_{2}}}\right)$$H_{*}\left(F_{x_{i-1}^{l_{2}}}\right)$$H_{*}\left(F_{\bigvee_{j\leq i+1}x_{j}}\right)$$H_{*}\left(F_{x_{i+1}^{l_{1}}}\right)$$H_{*}\left(F_{\bigvee_{j\leq i}x_{j}}\right)$$H_{*}\left(F_{x_{i-1}^{l_{1}}}\right)$$H_{*}\left(F_{x_{i}^{l_{1}}}\right),$$\scriptstyle{d}$$\scriptstyle{b}\scriptstyle{\sim}\scriptstyle{\sim}\scriptstyle{\sim}\scriptstyle{\sim}$$\scriptstyle{a}$$\scriptstyle{c}$ (25)
where $\bigvee$ denotes the coordinate-wise maximum, and the isometries are guaranteed by Equations 21 and 22. Such diamonds are called transposition diamonds, on which (52, Theorem 2.4) applies and states that if $\left\{\tau_{1},\dots,\tau_{N}\right\}$ is a reduced chain basis of the persistence module generated by applying the homology functor $H_{*}$ on the filtration $\left(K_{j}\right)_{1\leq j\leq N}$, then, there is an explicit updated basis $\left\{\mathrm{vine}_{i}(\sigma_{1}),\dots,\mathrm{vine}_{i}(\sigma_{N})\right\}$, that is reduced for the filtered chain complex $(\hat{K}_{j})_{1\leq j\leq N}$, where, given an index $j$, the complex $\hat{K}_{j}$ is defined as $\hat{K}_{j}:=\left<\sigma_{(i\,i+1)1},\dots,\sigma_{(i\,i+1)j}\right$>. This vine update follows a case study, that we follow below to show that the corresponding matching function is compatible.
  1. 1.
    The maps $a$ and $c$ are surjective of nullity 1, i.e., the added simplices $\sigma_{i}$ and $\sigma_{i+1}$ are negative. Let ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\in\left\{\tau_{1},\dots,\tau_{N}\right\}$ the chains generating these intervals, i.e., ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}=\partial\tau_{i}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}=\partial\tau_{i+1}$.
    1. (a)
      Assume that ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\in\ker(b)$. See Figure 14(a) for an illustration. In that case, (52, Theorem 2.4) guarantees that
      $\forall 1\leq j\leq N,\quad\tau_{j}\mapsto\mathrm{vine}_{i}\left(\tau_{j}\right):=\tau_{j}$ (26)
      is a reduced basis of the filtered chain complex $(\hat{K}_{j})_{1\leq j\leq N}$ and hence of the filtered chain complex $\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}$}. Hence, Equation 21 guarantees that for any index $j$, the matched bar endpoints, $\tau_{j}\mapsto\mathrm{vine}_{i}(\tau_{j})$, with filtration values $x_{j}^{l_{1}}$ and $x_{j}^{l_{2}}$ are not strictly comparable, and thus the induced matching function is compatible.
    2. (b)
      If ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\notin\ker(b)$, then (52, Theorem 2.4) guarantees that there exists $\alpha\in k\setminus\left\{0\right\}$ such that ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\in\ker(b)$. See Figure 14(b) for an illustration. In other words, in the quotient space $H_{*}(K_{i-1})$}, we have ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}=u_{1}\partial\sigma_{i}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}=v_{1}\partial\sigma_{i}+v_{2}\partial\sigma_{i+1}$ for some invertible constants $u_{1},v_{1},v_{2}\in k\setminus\left\{0\right\}$, with $\alpha u_{1}=-v_{1}$. Hence, we have:
      $\left\{\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}),\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})\right\}=\begin{cases}\left\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\right\}\quad\textnormal{if }{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})\leq{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}),\\ \left\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\right\}\quad\textnormal{otherwise,}\end{cases}$ (27)
      and the identity for the non-impacted chains:
      $\forall\tau_{j}\notin{\left\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v},\tau_{i},\tau_{i+1}\right\}},\quad\mathrm{vine}_{i}\left(\tau_{j}\right):=\tau_{j}.$ (28)
      The chains $\tau_{i}$ and $\tau_{i+1}$ are then updated such that the vine update commutes with the boundary, i.e., such that:
      $\mathrm{vine}_{i}(\partial\tau_{i})=\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})=\partial\mathrm{vine}_{i}(\tau_{i})\quad\textnormal{and}\quad\mathrm{vine}_{i}(\partial\tau_{i+1})=\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})=\partial\mathrm{vine}_{i}(\tau_{i+1}).$ (29)
      In particular, Equation 21 guarantees once again that (at least) all but four endpoint (two bars, with two endpoints each) matching are compatible. Furthermore, note that since the simplices’ order only differ from the permutation (ii+1)(i\,i+1) and the chains ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}$ belong to $K_{i-1}$, there exist two indices $j,k<i$ such that the inequality $x_{j}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})\leq{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})=x_{k}^{l_{1}}$ is equivalent to $x_{j}^{l_{2}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})\leq{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})=x_{k}^{l_{2}}$. We fix such indices $j$ and $k$.
      We follow the two different cases.
      1. i.
        In the first case, we have $\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}$ (with $\mathrm{vine}_{i}(\tau_{i})=\tau_{i}$) and $\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}$ (with $\mathrm{vine}_{i}(\tau_{i+1})=\tau_{i}+\alpha\tau_{i+1}$). Furthermore, since ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})\leq{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})$, we also have ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\right)={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\sigma_{k})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\right)=x_{k}^{l_{2}}$. Hence, Equation 21 guarantees once again that the birthpoint matching is compatible. The same goes for the deathpoints matching, since for $j\in\left\{i,i+1\right\}$, we match the endpoints $x_{j}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{j})$ and $x_{j}^{l_{2}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}(\tau_{j}))$, which are not strictly comparable. Thus, the matching function is compatible.
      2. $i$.
        In the second case, we have $\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}$ (with $\mathrm{vine}_{i}(\tau_{i})=\tau_{i}+\alpha\tau_{i+1}$) and $\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}$ (with $\mathrm{vine}_{i}(\tau_{i+1})=\tau_{i+1}$). In this case, we have ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})>{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})$, and we match the birthpoints
        $x_{j}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})\textnormal{ with }{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})=x_{j}^{l_{2}},\textnormal{ and,}$
        $x_{k}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})\textnormal{ with }x_{k}^{l_{2}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}),$
        and the deathpoints
        $x_{i}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{i})\textnormal{ with }{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}(\tau_{i}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{i}+\alpha\tau_{i+1})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\sigma_{i+1})=x_{i+1}^{l_{2}},\textnormal{ and }$
        $x_{i+1}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{i+1})\textnormal{ with }x_{i}^{l_{2}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}(\tau_{i+1}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{i+1})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\sigma_{i}).$
        Finally, the birthpoints are not strictly comparable hence compatible thanks to Equation 21 and the deathpoints thanks to Equation 23. The matched bars are thus compatible.
    ((a)) Illustration of case 1a.
    ((b)) Illustration of case 1b.
    Figure 14: Vineyard case: $\sigma_{i}$ and $\sigma_{i+1}$ negative.
  2. 2.
    The map $a$ and $c$ are injective of corank $1$, i.e., the simplices $\sigma_{i}$ and $\sigma_{i+1}$ are positive. See Figure 15(b) for an illustration.
    1. (a)
      Assume that ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}\in\mathrm{im}(d)$. Then, (52, Theorem 2.4) guarantees that the updated basis given by $\tau_{j}\mapsto\mathrm{vine}_{i}\left(\tau_{j}\right):=\tau_{j}$ is a reduced basis of $\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}$} as well. Using a similar argumentation as Case 1a, the matching function is compatible.
    2. (b)
      If there exists an $\alpha\in k\setminus\left\{0\right\}$ such that ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\in\mathrm{im}(d)$, i.e., ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\in H_{*}(\hat{K}_{i})$, i.e., in $H_{*}(K_{i-1})$, we have ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}=u_{1}\sigma_{i}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}=v_{1}\sigma_{i}+v_{2}\sigma_{i+1}$ , for invertible constants $u_{1},v_{1},v_{2}\in k\setminus\left\{0\right\}$ satisfying $\alpha u_{1}=-v_{1}$. We also have two cases
      $\left\{\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}),\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})\right\}=\begin{cases}\left\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}\right\}\quad\textnormal{if }{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\delta{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})\leq{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\delta{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}),\\ \left\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\right\}\quad\textnormal{otherwise,}\end{cases}$
      where ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\delta{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})$ (resp. ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\delta{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})$) is the first time in which a coboundary of $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u$} (resp. $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v$}) appears where, given a chain cycle $\tau$, we define ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\delta\tau):=\inf\left\{t\in l_{2}:\tau=0\textnormal{ in }H_{*}(F_{t})\right\}\in\mathbb{R}\cup\left\{+\infty\right\}$. When they exist, we consider the first coboundary $\tau_{j}$ (resp. $\tau_{k}$) in $\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}$} of $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u$} (resp. $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v$}) i.e., when the cycle $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u$} (resp. $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v$}) is not essential, we consider the index $j>i+1$ (resp. $k>i+1$) such that $\partial\tau_{j}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}$ (resp. $\partial\tau_{k}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}$), and hence ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\delta{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{j})$
      ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\delta{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{j})\quad\textnormal{ and respectively }\quad{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\delta{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{k})$ (30)
      Note that since the simplices’ order on $l_{1}$ and $l_{2}$ only differ from the permutation (ii+1)(i\,i+1), Equation 30 is also satisfied when the line $l_{1}$ replaced by $l_{2}$.
      The other simplices being updated by the identity i.e. by $\mathrm{vine}_{i}(\tau_{j}):=\tau_{j}$, unless their boundary is $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u$} or $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v$}, in which case there are updated as in previous Cases 1(b)i and 1(b)ii, by Equation 29.
      1. i.
        In the first case (${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\delta{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})\leq{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\delta{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})$), simlilarly to Case 1(b)i, we have $\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}$ (with $\mathrm{vine}_{i}(\tau_{j})=\tau_{j}$), and $\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}$ (with $\mathrm{vine}_{i}(\tau_{j+1})=\tau_{j}+\alpha\tau_{j+1}$); which ensures, since ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}\right)=x_{i+1}^{l_{2}}$ and ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})=x_{i+1}^{l_{1}}$, that this case also induces a compatible matching function. If $\tau_{j}$ exists, then ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}(\tau_{j}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{j})$ which induces a compatible endpoint matching. If $\tau_{j}$ and $\tau_{k}$ exist, then, as ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{j})<{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{l})$, we have ${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}(\tau_{k}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{j}+\alpha\tau_{k})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{k})$ which also induces a compatible endpoint matching. Hence, the vineyard barcode matching is compatible in this case.
      2. $i$.
        The second case (${\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\delta{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})<{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\delta{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})$) is similar to Case 1(b)ii, we pick $\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}$ and $\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}$. We match the birthpoints
        $x_{i}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})\textnormal{ with }{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}+\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})=x_{i+1}^{l_{2}},\textnormal{ and,}$
        $x_{i+1}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})\textnormal{ with }{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\sigma_{i})=x_{i}^{l_{2}}.$
        Equation 23 hence guarantees that these endpoints are strictly uncomparable, hence compatible. Furthermore, when $\tau_{k}$ exists, we match the deathpoints
        $x_{k}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{k})\textnormal{ with }{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}(\tau_{k}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{k})=x_{k}^{l_{2}}.$
        and when $\tau_{j}$ exist as well:
        $x_{j}^{l_{1}}={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{1}}}(\tau_{j})\textnormal{ with }{\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\mathrm{vine}_{i}(\tau_{j}))={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{j}+\alpha\tau_{k})={\left.\kern-1.2ptF\vphantom{\big|}\right|_{l_{2}}}(\tau_{j})=x_{j}^{l_{2}},$
        In both cases, these deathpoint matching are compatible thanks to Equation 21. The matching function is therefore also compatible.
    ((a)) Illustration of case 2a.
    ((b)) Illustration of case 2b.
    Figure 15: Vineyard case: $\sigma_{i}$ and $\sigma_{i+1}$ positive.
  3. 3.
    $a$ is injective of corank 1 and $c$ is surjerctive of nullity 1, i.e., $\sigma_{i}$ is positive and $\sigma_{i+1}$ is negative. See Figure 16(a) for an illustration. In this case, since these two chains are not of the same dimension, they do not interact together. Hence, we have $\mathrm{vine}_{i}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u}$ and $\mathrm{vine}_{i}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v})={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v}$ and Equation 21 guarantees that the matching function is compatible.
  4. 4.
    $a$ is surjective of nullity 1 and $c$ is injective of corank 1, i.e., $\sigma_{i}$ is negative and $\sigma_{i+1}$ is positive. See Figure 16(b) for an illustration. This case is symmetrical to Case 3.
    ((a)) Illustration of case 3.
    ((b)) Illustration of case 4.
    Figure 16: Vineyard case: $\sigma_{i}$ and $\sigma_{i+1}$ of different signs.
Remark 12.
Note that, in this proof, and in the specific cases for which the vineyards algorithm makes a matching choice (which are Cases 2b and 1b), these corresponding choices are made so that one cycle is left unchanged. Furthermore, as permuted simplices are guaranteed to be strictly incomparable (Equation 23), both choices (i.e., permuting $i$ and $i+1$ or not) induce a compatible matching. Also notice that, in the generic case, this case study allows to span the set of all possible matchings between two lines whose induced orders only differ from a single (ii+1)(i\,i+1) transposition.
Remark 13 (Extension to free presentations).
In this proof, we only used the framework of simplicial complexes in order to ensure that simplices can be added one by one, which in turn guarantees that the corresponding linear maps are either injective of corank 1, or surjective of nullity 1. However, given a multi-parameter persistence module $\mathbb{M}$, such an assumption on linear maps can also be guaranteed using a finite free presentation of $\mathbb{M}$ (see (13, Section 7)). Hence, computing a minimal presentation can be seen as a pre-processing step of our MMA algorithm.
Remark 14 (Coxeter decompositions).
Given a permutation of simplices between two lines, one might wonder how to decompose it into a product of transpositions (ii+1)(i\,i+1) in order to apply the vineyards algorithm. This can be achieved in the three following steps. A permutation $\sigma\in\mathfrak{S}_{n}$ can always be decomposed into a product of cyclic permutations, i.e., permutations $\sigma_{i_{1},\dots,i_{k}}\in\mathfrak{S}_{n}$ satisfying:
[…]
Then, using the fact that any cyclic permutation $\sigma_{i_{1},\dots,i_{k}}=(i_{1}\,i_{2})\circ(i_{2}\,i_{3})\circ\cdots(i_{k-1}\,i_{k})$ is a product of transpositions, it follows that every permutation is a product of transpositions as well. Finally, if $(i,j)$ is a transposition with $1\leq i<j\leq n$, using the following relation:
$(i\,j)=(i\,i+1)\circ\cdots\circ(j-1\,j)\circ(j-2\,j-1)\circ\cdots\circ(i\,i+1),$
one can see that every permutation is in fact the product of adjacent transpositions. In practice, this product can be retrieved using a sorting algorithm, such as the insertion sort, or the bubble sort (see, e.g., (42)).
Remark 15 (Lazy vineyard update).
Note that the (ii+1)(i\,i+1) swaps can be directly inferred with the approach provided in (38, Section 4, Lazy minimazation). More formally, using the same notations, the swaps (ii+1)(i\,i+1) only occur precisely when two incomparable filtration values satisfy $x_{i}^{l}=x_{i+1}^{l}$ for some line $l$, which, in the two-parameter case, corresponds to the presence of a line crossing the point $(\max\left\{(x_{i})_{1},(x_{i+1})_{1}\right\},\max\left\{(x_{i})_{2},(x_{i+1})_{2}\right\})$. This suggests that our MMA algorithm can be trivially extended to fibered barcodes involving only such lines (instead of $\delta$-grids of lines), thus reducing its running time. We stick to grids in this article for the sake of clarity.
Proposition 17 (Vineyards is compatible for $n=2$).
Let $\mathbb{M}$ be a $2$-parameter persistence module, and $L$ be an ordered set of diagonal lines $L:=(l_{i})_{1\leq i\leq N}$ with increasing basepoints. [^6] Ordering $L$ in such a way is possible precisely because $n=2$. For each index $1\leq i\leq N$, let $\sigma_{i}$ be an arbitrary compatible matching function between ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{i}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{i+1}}}\right)$ (obtained with, e.g., Proposition 16). Then, the matching function $\sigma=(\sigma_{i})_{1\leq i\leq N}$ is a compatible matching function on $L$.
Proof.
First, note that, given a compatible matching function $\sigma$ between two diagonal lines $l,l^{\prime}\in L$, and letting $b\in{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),b^{\prime}\in{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right)$ be any non-trivial pair of matched bars, then, assuming that, e.g., $(x,0)\leq(x^{\prime},0)$ are the basepoints of $l$ and $l^{\prime}$ respectively, one has $\min(b)_{1}\leq\min(b^{\prime})_{1}$ and $\max(b)_{2}\geq\max(b^{\prime})_{2}$. See Figure 17.
Figure 17: Let $l_{1}<l_{2}$ be two diagonal lines of $\mathbb{R}^{2}$, and $x$, $y$ be two matched points in $l_{1}$ and $l_{2}$ respectively. Then, if the matching is compatible, one must have $x_{1}\leq y_{1}$ and $x_{2}\geq y_{2}$.
Now, let $x\in l_{1}$ be the grade of a given bar endpoint. A point $y$ is strictly incomparable with $x$ if either:
  1. 1.
    $y_{1}\leq x_{1}$ and $y_{2}\geq x_{2}$, or
  2. 2.
    $y_{1}\geq x_{1}$ and $y_{2}\leq x_{2}$.
Now, assuming that $y\in l_{2}$ and $x$ and $y$ are strictly incomparable, Case 1 can be excluded. Indeed, in that case, there exist constants $x_{0},y_{0},s,t\in\mathbb{R}$ such that $x=(x_{0}+s,s)$ and $y=(y_{0}+t,t)$, and hence:
[…]
This shows that two bar endpoints matched by a compatible matching function between two consecutive diagonal lines with increasing basepoints must satisfy Case 2, and thus, that a sequence of compatible matching functions is still compatible. ∎
7 Numerical experiments
In this section, we showcase the performances of MMA. More precisely, we empirically study how the output quality and running time depend on the precision $\delta$ and the number of lines $|L|$. Then, we compare the running times of MMA with those of Rivet (47) and the elder-rule staircode (ERS) (17), which are our closest competitors in terms of producing visual and interpretable descriptors of persistence modules. Finally, we investigate how running time is affected by the number of filtrations. All experiments were done on a laptop with AMD Ryzen 4800 CPU and 16GB of RAM. Our code is part of the multipers library (50) and is publicly available at https://github.com/DavidLapous/multipers.
Interpretation.
We first qualitatively show how to interpret our candidate decompositions on a toy dataset in Figure 18, using two different two-parameter filtrations (Čech and (edge-collapsed) Rips with sublevel sets of codensity) on a point cloud, computed using (2). In these examples, we consider a coordinate $x\in\mathbb{R}^{2}$ that is included in the support of some specific summands of the candidate decomposition, i.e., in some specific colored shapes in the plot. Then, we look back at the filtration at this coordinate $x$ and we identify the corresponding cycles. Note that, at each of these points $x$, the cycle representatives (from Theorem 15) are already calculated when running MMA and can thus be used for interpretation without additional computational cost.
((a)) (Left) The Čech complex, at radius $0.2$, of the points with codensity values larger than $3$. (Right) The corresponding candidate decomposition, with a red dot at the coordinates fixed by the radius and codensity values used on the left.
((b)) (Left) The (edge-collapsed) Rips complex, at radius $0.4$, of the points with codensity values larger than $3$. (Right) The corresponding candidate decomposition, with a red dot at the coordinates fixed by the radius and codensity values used on the left.
Figure 18: Interpretation of candidate decompositions computed by MMA on a toy dataset.
We also provide another example, containing $25,000$ points uniformly sampled on the unit square (noise), and $25,000$ points sampled on three distinct annuli with different sizes concentration levels (signal). See Figure 19. As before, we then compute the Čech-codensity filtration on this point cloud. The first cycle is very dense, so it should appear quickly w.r.t. the codensity filtration (i.e., the lower part of the two-parameter filtration), and it is also small, so it is expected to die quickly w.r.t. the Čech radius parameter. The second cycle is bigger, and slightly less dense, so we can expect it to appear later and survive more on the right side, and the same goes for the third one. The fourth one appears when the Čech radius is large enough (in order to connect the three previous cycles together), and the condensity parameter is large enough as well (such that the three previous cycles are visible).
Figure 19: (Left) The point cloud dataset, colored with the (estimated) density of the sampling. (Right) The candidate decomposition produced by MMA. One can see that four interval summands clearly stand out, and can be interpreted as described in the text. The summands that are induced by noise are all located on the rainbow strip on the left side.
Data sets and filtrations.
In our next two experiments, we focus on two real-world data sets of point clouds. The first ones, called LargeHypoxicRegion, were obtained from immunohistochemistry in (58). These are made of a few thousand points, each representing a single cell. The others were obtained from applying time-delay embedding in $\mathbb{R}^{2}$ on time series taken from a few data sets (Wine, Plane, OliveOil, Coffee) from the UCR archive (25). On all of these data sets, we computed bi-filtrations using the standard Vietoris-Rips filtration, and the superlevel sets of a Gaussian kernel density estimation (with bandwidth parameter $0.1d$ where $d$ is the diameter of the dataset), and we applied MMA and its competitors on the corresponding multi-parameter persistence modules in homology dimensions 0 and $1$ (note that the ERS can only be computed in degree 0). A typical example of an MMA representation is given in Figure 20.
In our third experiment, we measure the dependence on the number of filtrations using a synthetic data set obtained by sampling $300$ points in the unit square $0,1]^{2}$, computing their Alpha simplicial complex, and generating $n$ random filtration values on the points.
Finally, it is worth noting that we used multi-parameter edge collapses (3) in order to simplify the multi-parameter persistence modules (without losing information) as much as possible before applying Rivet and MMA. The timing of this simplifications are not taken into account, but they are not the computational bottleneck of our computations. Furthermore, for the LargeHypoxicRegion, we thresholded the Rips edges at $0.02$, leading to simplicial complexes of $\sim\!85k$ and $\sim\!125k$ simplices respectively, after simplifications.
Figure 20: (left) LargeHypoxicRegion2, colored with a kernel density estimation of bandwidth $0.01$, (middle, resp. right) Interval decomposition of degree 0 (resp. 1) homology given by MMA. There are $\approx 28k$ non-trivial intervals, each one having a unique color.
Convergence of MMA.
We first empirically validate Propositions 3, 4 and 5 by measuring how far is the output of MMA from the data when the precision parameter $\delta$ decreases and the number of lines in $L$ increases. For this, we used the bottleneck distances between the fibered barcodes (on $100$ random diagonal lines) of the outputs of MMA and the ones of the underlying modules as a proxy for the interleaving distances (since they are practically very difficult to evaluate). We call this the estimated matching distance. Results are displayed in Figure 21. One can see that the convergence is empirically linear with the number of lines $|L|$, which is in line with Propositions 3, 4 and 5 (since $|L|$ increases linearly as $\delta$ decreases for a fixed $n$). Note that the distance even reaches 0 on a few cases.
Figure 21: Convergence on UCR data sets (left) and immunohistochemistry data sets (right).
Running times.
We now compare the running times of MMA with those of Rivet and the ERS. Results can be found in Table 1. It is worth noting that on several occasions, Rivet and the ERS could not produce outputs in reasonable time, due (among other things) to large memory consumption. On the other hands, the fact that MMA produces discretely presented intervals allows to encode them in a sparse manner with their corners. Note that computations with Rivet can be also approximated by coarsening the filtration values, and thus the module. In practice, this corresponds to restricting the 2-module to a $\kappa=\kappa_{x}\times\kappa_{y}$ grid, where $\kappa_{x},\kappa_{y}\in\mathbb{N}$ are the resolutions in both axes, see (47, Section 8.2). The parameters $\kappa$ and $|L|$ have (roughly) the same role, and can be related with $\sqrt{\kappa}\simeq|L|$ (for a given, prescribed precision). Overall, we find that the running times of MMA outperforms its competitors, except when $\kappa$ is very small. However, in this case, the corresponding output of Rivet is very crude as the module is restricted on just a few points, whereas MMA produces intervals that are accurate along whole straight lines: we find that for large data sets, MMA is the only method that can produce accurate representations.
Rivet ERS MMA
$\kappa=10^{2}$ $\kappa=50^{2}$ $\kappa=100^{2}$ $|L|=100$ $|L|=1000$ $|L|=10,000$
Coffee $0.21\pm 0.01,0.18\pm 0.01$ $9.75\pm 5.92,0.35\pm 0.12$ $0.95\pm 0.56$ $34\pm 0.04 0.0$$093\pm 0.001 0.0$$024\pm 0.001$ $0.16\pm 0.008$
Plane $0.19\pm 0.005,0.18\pm 0.03$ $4.36\pm 2.24,0.28\pm 0.04$ $33.3\pm 17.5,0.56\pm 0.17$ $09\pm 0.03 0.0$$04\pm 0.0 0.0$$12\pm 0.001 0.0$$095\pm 0.004$
Wine $0.21\pm 0.003,0.19\pm 0.007$ $8.50\pm 2.00,0.22\pm 0.01$ $8\pm 0.023 0.22±0$$2\pm 0.04$ $004\pm 0.0 0.0$$016\pm 0.0$ $0.129\pm 0.002$
OliveOil $0.21\pm 0.004,0.19\pm 0.002$ $5.55\pm 1.20,0.31\pm 0.016$ $0.82\pm 0.17$ $1.39\pm 0.03$ $026\pm 0.001 0.0$$058\pm 0.001$ $0.37\pm 0.006$
Worms $0.29\pm 0.082,0.23\pm 0.23$ $19.9\pm 14.4,4.60\pm 5.0$ $31.36\pm 36.24$ $3.85\pm 0.1$ $0.22\pm 0.11$ $0.34\pm 0.15$ $1.35\pm 0.41$
LargeHypoxicRegion1 $1.73,2.88$ ,234--,234 ,--,-- -- $26.4$ $26.6$ $59.4$
LargeHypoxicRegion2 $2.39,6.04$ ,--,-- ,--,-- -- $57.3$ $54.3$ $102.9$
Table 1: Mean and variances of the running times (s) for Rivet, the ERS and MMA. We provide both degree 0 (left) and 1 (right) homology timings for Rivet, whereas the timings of MMA include both. The double dashes correspond to out of memory errors, i.e., a memory usage that is over $12\mathrm{GB}$.
Interestingly, computing 0-dimensional homology is sometimes slower than 1-dimensional homology for Rivet; as it relies on computing minimal free presentations, we think that this comes from the fact that minimal presentations in homology dimension 0 can be more complex than their counterparts in homology dimension 1, i.e., they have much more generators. We also investigate how running times of MMA depend on the number of lines. Unsurprisingly, one can see from Figure 22 that running time increases with the number of lines. However, the dependency looks empirically sublinear, which could come from the fact that even though there are more lines, these lines are closer to each other, and thus matching them with vineyards requires fewer computation steps. This is also highlighted by the running times of LargeHypoxicRegion2, Table 1 which are smaller when computing it over 10001~000 lines than $100$ lines.
Figure 22: Running time (s) needed to run MMA on the UCR datasets.
Figure 23: Running time (s) of MMA w.r.t. the number $n$ of filtrations.
Dependence on number of filtrations.
Finally, we investigate how the running times of MMA depend on the number $n$ of filtrations. Although most of the approaches in the literature are limited to $n=2$ parameters, one can see from Figure 23 that MMA can produce outputs in a reasonable amount of time for up to $n\simeq 10$ parameter filtrations. As expected from the complexity of MMA in Equation (9), the running times increase exponentially with $n$.
8 Conclusion
In this article, we present MMA: a new algorithm for computing topological descriptors for multi-parameter persistence modules taking the form of candidate decompositions. Our algorithm has complexity and running time that are controlled by user-defined parameters, and enjoy several approximation and stability properties. We also showcased the performances of MMA on synthetic and real data sets. Our code is part of the multipers library (50) and is publicly available at https://github.com/DavidLapous/multipers.
Along the way, we identified several directions for future work.
  1. 1.
    While the outputs of MMA satisfy approximation guarantees when computed with compatible matching functions, they remain arbitrary to some extent, as discussed at the end of related work and Figure 8.
    We hypothesize that, in the general case, the family $\mathcal{F}$ of approximate decompositions obtained from MMA by varying $\sigma$ across an appropriate family of compatible matching functions $\Sigma$, i.e., $\mathcal{F}=\{\texttt{MMA}{}({\mathbb{M}},L,\sigma)\}_{\sigma\in\Sigma,\delta>0}$, is a complete topological invariant of the module $\mathbb{M}$.
  2. 2.
    Practically speaking, the existence and construction of compatible matching functions for $n$-parameter persistence modules with $n>2$ is an open question. We hypothesize that matching functions computed from tracking representative cycles, in a similar way than the construction of the graphcode (39, 40), could provide a step towards that direction. Another possibility includes designing a convex optimization problem that would converge to compatible (and potentially stable) matching functions.
Statements and Declarations
Ethical Approval
Not applicable.
Competing interests
No competing interests.
Authors’ contributions
D.L. M.C. and A.B worked out the proofs and wrote the manuscript, D.L. did the numerical experiments. All authors reviewed the manuscript.
Funding
D.L. was funded by the French government through the 3IA Côte d’Azur Investments, ANR-19-P3IA-0002. M.C. was supported by Agence Nationale de la Recherche through ANR JCJC TopModel ANR-23-CE23-0014, and by the French government, through the 3IA Cote d’Azur Investments in the project managed by the National Research Agency (ANR) with the reference number ANR-23-IACL-0001.
Availability of data and materials.
All of the non-synthetic data sets used are publicly available, and listed below.
  1. 1.
    The immunohistochemistry datasets can be retrieved from
    https://github.com/MultiparameterTDAHistology/SpatialPatterningOfImmuneCells, and
  2. 2.
    The time series datasets can be retrieved from http://www.timeseriesclassification.com/dataset.php.
All of the code used is publicly available as well.
  1. 1.
    Our code, in multipers, at https://github.com/DavidLapous/multipers,
  2. 2.
  3. 3.